### v1

# Stochastic gravitational-wave background search using data from five pulsar timing arrays

Abstract
Using public data from five pulsar timing arrays (PTAs), we search for a stationary, isotropic, and unpolarized nHz stochastic gravitational-wave background (SGWB). We use pulse-time-of-arrival data from 121 pulsars, which is more informative than previous searches, carried out separately by the individual PTA collaborations using only their own data. For pulsars observed by more than one PTA, we employ a new “direct combination” method to merge their astrophysical models and data. This avoids the challenge of reconciling the different PTA timing models to obtain a single “best” model. A central result of our analysis is posterior probability distributions for the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency. While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance—assessed via a Bayesian odds ratio and through noise-marginalized false-alarm probabilities ($p$-values) for three different detection statistics—remains below the conventional $5\sigma$ threshold for a confident detection. We also reconstruct the inter-pulsar timing-residual correlation, as a function of the angle $\theta$ between the pulsar lines of sight. This is consistent with the curve predicted by Hellings and Downs (HD).
Introduction – Pulsar timing arrays (PTAs) aim to detect low frequency gravitational waves (GWs), by searching for the effects that such waves induce on the arrival times of pulsar pulses. These GWs create “timing residuals” which are correlated among different pulsars [1, 2]. Decades-long datasets with submicrosecond timing precision yield GW amplitude (strain) sensitivities of order $A_{gw}\sim 10^{-15}$ at nHz frequencies.
Orbiting pairs of supermassive black holes (SMBHs) are one possible source of GWs. If the Universe contains many such pairs, then the incoherent superposition of their GW emissions would create a stochastic GW background (SGWB) [3, 4, 5] with a power-law spectrum in the nHz band.
Different PTA collaborations have reported varying levels of evidence for a nHz SGWB. The Parkes PTA (PPTA [6]) observed 30 pulsars over 18 years, concluding that the data showed “no support for or against”. The European and Indian PTAs (EPTA and InPTA [7]) observed 25 pulsars, and reported “marginal evidence” in the 25-year dataset and “evidence” when only the last 10 years of data were analyzed. The Chinese PTA (CPTA [8]) observed 57 pulsars over three years, and reported “some evidence”. The North American Nanohertz Observatory for Gravitational Waves (NG [9]) observed 67 pulsars over 15 years and reported “compelling evidence”. Most recently, the MeerKAT PTA (MPTA [10]) observed 83 pulsars over 4.5 years, and, depending on assumptions about pulsar noise processes, reported either “high significance” or “no significance” (also see [11]). None of these pass the threshold of $5\sigma$ significance required to claim detection [12].
Figure 1: A Mollweide projection showing sky locations of the 121 pulsars included in our analysis. Colors indicate how many PTAs contributed data for each pulsar.
All of these PTAs (apart from CPTA) have released their data to the public. Using this full public dataset increases the total number of distinct pulsars to $N=121$, as shown in Fig. 1. PTAs search for an SGWB by correlating pairs of pulsars, whose number scales as $N^{2}$. Furthermore, since most pulsars are observed by more than one collaboration, there are also more data samples per pulsar. Hence, analysis of the full dataset should significantly increase the GW strain sensitivity and the detection confidence [13, 14, 15, 2].
We analyze this five-PTA dataset [16, 17, 18, 19, 20] using a recently proposed “direct combination” method [21, 22]. This allows us to combine the data with only minor modifications, as described below. Tests of this method, carried out with the International Pulsar Timing Array (IPTA) Data Release 2 (DR2) combined data set from 2019 [14], suggest that it produces results that are identical to traditional (but more time-consuming) data combination techniques.
Aside from this, our analysis is standard. We use the same tools and methods as the individual PTAs, as implemented in Enterprise [23] and in the next-generation PTA data-analysis package Discovery [24]. Both are based on parameterized likelihood models and employ Markov chain Monte Carlo methods. These produce posterior samples (here called “posteriors”) that approximate the inferred posterior probability distributions of the model parameters.
Methods, Models, and Conventions – Because we use standard methods, this paper is brief. When we follow the PTA analyses described above [6, 7, 8, 9, 10], we give only a brief description and cite the relevant literature. If we do something different, then more details are provided.
As is usual in this field, our analysis considers two alternative hypotheses or “models”: – The signal hypothesis, denoted by “HD”, is a model with two contributions to the pulsar pulse arrival-time residuals. The first is pulsar noise, assumed to be (a) uncorrelated between pulsars and (b) a sum of stationary “red” (RN) and “dispersion measure” (DM) noise plus a nonstationary “white noise” (WN) component. The second contribution arises from a stationary, unpolarized and isotropic SGWB, as described by the general theory of relativity (GR). This SGWB component, determined by parameters $A_{gw}$ and $\gamma_{gw}$ defined below, induces timing-residual correlations between pulsar pairs, which are assumed to follow the Hellings and Downs (HD) prediction [25, 26, 2] given in (2). Both contributions are assumed to arise from (the sum of) many independent sources; by the central limit theorem, the resulting timing residuals form a zero-mean Gaussian process, which is fully characterized by its covariance matrix $\bf C$.
– The null hypothesis is called “common uncorrelated red noise” and is denoted by CURN. Again, the timing residuals are assumed to be drawn from a zero-mean Gaussian process, but this time entirely characterized by the covariance matrix ${\bf C_{0}}={\rm diag}\,\bf C$. Since our null hypothesis is defined via our signal hypothesis, it has exactly the same parameters/noise-models as the signal hypothesis. The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the HD correlations [27]. It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms. Thus, CURN does not correspond to the absence of GWs, but rather to “GWs” which create arrival-time fluctuations lacking the inter-pulsar correlations predicted by GR. Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2]. CURN is considered a more conservative null hypothesis than that of “no GWs”.
Using Discovery [24] as a tool, and conditioning the Gaussian likelihood on the data and the priors, we generate posteriors for the HD and CURN models. These are sets of parameter values that are most the consistent with the data and priors; they characterize the joint posterior probability distributions of the model parameters.
For this work, the most important model parameters are the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the SGWB power-law power spectrum. To define these, let ${\rm{d}}\rho_{gw}$ denote the SGWB energy density in the frequency range $f,f+{\rm{d}}f]$ and let $\rho_{cr}=3c^{2}H_{0}^{2}/8\pi G\simeq 7.7\times 10^{-10}\ \mathrm{J}/\mathrm{m}^{3}$ denote the critical (closure) energy density of the Universe. Here, $c$ is the speed of light, $G$ is Newton’s gravitational constant, and $H_{0}\simeq 67.4~\mathrm{km}/\mathrm{s\,Mpc}\simeq 2.2\times 10^{-18}\,\mathrm{Hz}$ is the current Hubble expansion rate. Then
$\Omega_{gw}(f)\equiv\frac{f}{\rho_{cr}}\frac{{\rm{d}}\rho_{gw}}{{\rm{d}}f}=\frac{2\pi^{2}}{3H_{0}^{2}}A^{2}_{gw}f_{r}^{2}\Bigl(\frac{f}{f_{r}}\Bigr)^{-\gamma_{gw}+5}\,$, (1)
where $f_{r}=1/\mathrm{yr}\simeq 31.7\,\mathrm{nHz}$ is a standard reference frequency. Together with our previous assumptions, this spectral model (assumed to apply in the range $2<f/\mathrm{nHz}<200$) fully characterizes the SGWB.
One result of this analysis is posterior probability distributions for $A_{gw}$ and $\gamma_{gw}$. We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to take the value $13/3$. The latter would be expected from a population of SMBHs in circular binaries, provided that GWs dominate the energy loss at nHz frequencies.
One way to assess detection confidence is via frequentist detection statistics [27]. We present results for three different statistics that are used within the PTA community. All three statistics $D$ are quadratic functions of the timing residuals, with weights determined by $\bf C$ and $\bf C_{0}$.
Each of the posteriors corresponds to a different model universe, with its own amplitudes and exponents for the different noise sources, and thus with its own covariance matrices $\bf C$ and $\bf C_{0}$. This means that the values of the detection statistics $D$, and the probability ($p$-value) of obtaining them under CURN, differ from one posterior sample to the next. Thus, to assess detection confidence, we present the posterior distributions of these $p$-values. In the literature, its mean is called the “noise-marginalized false-alarm probability” [28] or “posterior predictive $p$-value” [29].
Some past work uses ad-hoc methods for assessing the false-alarm probability $p$. “Sky scrambling” and “phase shifting” were intended to “remove any signal from the data”. The first leaves the data invariant but assigns a large number of random sky positions to the different pulsars, thus destroying the HD correlation [30, 31]. The second retains the pulsar’s positions, but randomizes the phases of the (frequency-domain Fourier amplitudes of the) data [31]. However, neither of these methods properly reproduces the null distribution of $D$ [32]. For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
We determine the false-alarm probability $p$ using a correct method [33, 34, 32, 27]. Since $D$ is a quadratic form in the data, its distribution under CURN is a generalized $\chi^{2}$ distribution [33, 32, 27]; the probability that it exceeds the observed value in one of our posterior samples is found analytically.
To aid interpretation, we also report the $p$-values as “equivalent $q\sigma$” levels: $q>0$ is chosen so that a random draw from a zero-mean Gaussian with variance $\sigma^{2}$ has probability $p$ of exceeding $q\sigma$. This “one-sided $p$ value” satisfies $2p=\mathrm{erfc}(q/\sqrt{2})$. For example, $1\sigma$ corresponds to $p=16\%$, $2\sigma$ corresponds to $p=2.3\%$, $3\sigma$ corresponds to $p=0.14\%$, $4\sigma$ corresponds to $p=3.2\times 10^{-5}$, and $5\sigma$ corresponds to $p=2.9\times 10^{-7}$. The latter, a false-alarm probability of less than one per three million, is the traditional threshold for confident detection [12].
Analysis of the five-PTA dataset – The data from the five PTAs come from 121 millisecond pulsars, as shown in Fig. 1. These consist of the second EPTA data release (10-year subset) [35, 16], the first InPTA data release [36, 17], the MPTA 4.5-yr data release [37, 18], the NG 15yr data release [38, 19], and the third PPTA data release [39, 20]. Each PTA’s data consists of one .par file and one or more .tim files, per pulsar.
We merge these using the direct combination method [21, 22]. The .tim files, containing pulse times-of-arrival (TOAs) and their uncertainties, are used “as is”. The .par files, containing the astrophysical models, and some detector-specific parameters, are modified (in effect) as follows, one pulsar at a time. (This is also the first step of “manual” data combination.)
  1. 1.
    For pulsars observed by more than one PTA, a “reference PTA” is selected via the arbitrary ordering $\rm{NG}>\rm{EPTA}>\rm{PPTA}>\rm{MPTA}$. (InPTA is not listed because its pulsars are observed by other PTAs.) Any other PTAs that have data for this pulsar are called “target PTAs”.
  2. 2.
    Since some pulsars have both J- and B-style names, the name is copied from the reference .par file to the target .par files.
  3. 3.
    Any .par file that uses the TCB time standard is converted to TDB using Tempo2 [40] tools.
  4. 4.
    If the reference .par file has a DMX model for DM fluctuations (NG and InPTA only), then it is replaced by a Gaussian process (DMGP) model. The parameter DM0 is initialized to the mean DMX value, and DM1 and DM2 are initialized to zero.
  5. 5.
    If the reference .par file has a DMGP model, then the DM0 value is retained, and DM1 and DM2 are initialized to zero.
  6. 6.
    The astrophysical model $\beta$ [1, Eq. 7.2], including sky position, proper motion, spin frequency, spin-down parameters, binary orbital parameters, and DM values, is copied from the reference .par file to the target .par files.
Detector-specific parameters in the .par files, for example, “JUMP” for time offsets and “FDx” for narrow-band frequency offsets, are unchanged. This allows for template model shape differences between PTAs.
For four pulsars, ad-hoc “tricks” are used to maintain phase connection between the .tim file TOAs. For the binary pulsar PSR J2241-5236 (MPTA, PPTA), we retain the MPTA 11-parameter model and the PPTA 18-parameter model for the orbital frequency evolution. The other parameters follow 6 above. For PSR J0610-2100 (MPTA, NG) and PSR J1024-0719 (EPTA, MPTA, NG, PPTA), we keep the DMX model for the NG data and the DMGP model for the remaining PTAs. For PSR J1804-2858 (MPTA), we retain the values of DM1 and DM2 rather than initializing them to zero.
The direct combination method [21, 22] works because the only significant change it makes is to the parameters $\beta$, item 6 above. While their exact values are unknown, every pulsar has unique values for these parameters, for example, the sky position. Those enter the analysis via a nominal value $\beta_{0}$ [1, Eq. 7.2] and via a matrix $\mathbf{M}$} [1, Eq. 7.4], which is the rate of change of the pulse arrival-time with respect to variations $\epsilon$ in $\beta$. In the (tiny) region of interest, $\mathbf{M}$} varies slowly with those parameters, so computing $\mathbf{M}$} with the same $\beta_{0}$ for a given pulsar in all PTAs introduces negligible errors. The subsequent marginalization over $\epsilon=\beta-\beta_{0}$ [1, Eq. 7.6] fully accounts for the uncertainties in $\beta$ and for the variations in its nominal value across PTAs. Similarly, the RN and DM noise are intrinsic to the pulsar and to the interstellar medium (ISM), so must be the same for all PTAs.
The next analysis step is single-pulsar noise modeling. The goal is to determine each pulsar’s WN covariance matrix $\mathbf{N}$}, which appears in the Gaussian likelihood [1, Eq. 7.27]. This characterizes the timing residuals fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eq. 7.2,7.26]. Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
Single-pulsar noise modeling determines two or three parameters per pulsar and backend. The WN covariance matrix $\mathbf{N}$} is set by these parameters and by the .tim file contents. NG, PPTA, MPTA, and seven EPTA backends use the parameters $EFAC$, EQUADEQUAD, and $ECORR$ [38, Eq. 2]. InPTA and the remaining EPTA backends only use $EFAC$ and $EQUAD$ [41].
Parameters Uniform on range Stage
EFAC (0.01, 10) S
$\log_{10}$ EQUAD (-8.5, -5) S
$\log_{10}$ ECORR (-8.5, -5) S
$\gamma$ (1,7) S, F
$\log_{10}A$ (-20, -11) S, F
Table 1: Priors for the WN parameters, and for the amplitudes/exponents of the power-law RN, DM and SGWB noise spectra. “S” denotes single pulsar, and “F” denotes full PTA.
The WN parameters for a given pulsar are obtained from all of its TOA data (meaning, all .tim files). These are analyzed using JAXopt [42], with priors taken from the Table 1 rows with an “S” in the final column; the amplitude and exponent priors are used for all power laws. For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model; for $\gamma>1$, the $f\to\infty$ divergences are absent [43, 44]. Note that $EQUAD$ has the units used in Tempo [45], Tempo2 [40] and PINT [46], and is often written as $T2EQUAD$. The output is a “WN dictionary” of best-fit $EFAC$, EQUADEQUAD and (potentially) $ECORR$ values, which are used in subsequent analysis to obtain $\mathbf{N}$} [38, Eq. 2].
For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA-analysis values. This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more reliable estimates of the WN parameters.
We drop TOAs for any backends which have $EFAC<0.1$ or $EFAC>10$, since this indicates that the backend or template matching was unstable, or that RFI or calibration or instrumental issues were present. This removes $0.17\%$ of the TOAs. Our final data for analysis comes from 25 EPTA pulsars, 13 (of 14) InPTA pulsars, 83 MPTA pulsars, 68 NG pulsars, and 31 (of 32) PPTA pulsars. (Data from InPTA pulsar J1022+1001 are dropped because all the $EFAC$s are larger than 10. Data from PPTA pulsar PSR J1741+1351 are dropped because there were only 16 observations.)
The final stage is multi-pulsar analysis, following the procedures described in [1, 9]. We use both Enterprise [23] and the next generation PTA data-analysis package, Discovery, which is optimized via JAX for graphics processing units (GPU) [47]. The use of Discovery has significantly sped up our analysis.
Within the Discovery framework, we use the priors listed in Tab. 1 with an “F” in the final column. As described above, we consider two priors for the SGWB power-law power-spectrum exponent $\gamma_{gw}$. One is the uniform range given in Tab. 1, the other is fixed to $\gamma_{gw}=13/3$. Our analysis uses the Hamiltonian No-U-Turn Sampler (NUTS) [48, 49], as implemented in NumPyro [50, 51, 52]. NUTS is the default for Discovery [24], which provides the log-likelihood gradients needed by Hamiltonian samplers. It is much faster than the PTMCMC sampler [53] used by the individual PTAs when they analyzed their own data.
We generated $20{\kern 0.83298pt}480$ posterior samples under the HD signal hypothesis, and the same number of posteriors under the CURN hypothesis. Each gives the amplitudes and exponents of the power-law models describing the pulsar (red and DM) noise, and the SGWB. We use these to evaluate the frequentist detection statistics and their corresponding false-alarm probabilities ($p$-values) for each posterior sample.
Figure 2: The posteriors for the amplitude and exponent of the SGWB power spectrum, as defined in (1).
Results of the five-PTA analysis – One important result of this analysis are the posterior distributions of the SGWB power-spectrum amplitude $A_{gw}$ and exponent $\gamma_{gw}$ as defined in (1). These are shown in Fig. 2. The amplitude posterior is sharply peaked away from zero, which is consistent with (or evidence for) the presence of an SGWB. As expected, the posteriors are very similar for the two hypotheses. This is because, if there are $\lesssim 1500$ pulsars (see [54, Fig. 10]), most of the SGWB evidence arises from autocorrelations rather than cross-correlations—–and by construction, CURN includes these autocorrelations.
Model $\log_{10}A_{gw}$ $\gamma_{gw}$
CURN with free $\gamma_{gw}\qquad$ $14.24^{+0.11}_{-0.11}\qquad$ $3.30^{+0.28}_{-0.27}$
CURN with $\gamma_{gw}=13/3\qquad$ $14.69^{+0.04}_{-0.04}\qquad$
HD with free $\gamma_{gw}\qquad$ $14.28^{+0.12}_{-0.12}\qquad$ $3.32^{+0.28}_{-0.28}$
HD with $\gamma_{gw}=13/3\qquad$ $14.72^{+0.05}_{-0.05}\qquad$
Table 2: Median posteriors values for $A_{gw}$ and $\gamma_{gw}$, and their 68% containment intervals (gray bands in Fig. 3).
Happily, these posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6]. They are more constraining than the posteriors obtained by the individual PTAs: the uncertainty ellipses in the bottom-left plot of Fig. 2 are smaller than those obtained by intersecting the individual PTA posteriors [55]. For example, the $\text{area}=\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse (for the HD hypothesis) is 45% of that shown in [55, Fig. 1, right]. As explained earlier, this is as expected for an SGWB: joint analysis yields more evidence and is more constraining than naive intersection of the separate PTA results.
To claim an SGWB detection, there are several hurdles to overcome [12]. One must show that the data is consistent with the HD signal hypothesis, and that it is very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis. Clearly, the estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are almost identical under both hypotheses.
One way to assess significance is through null-hypothesis testing [56] using a frequentist statistic. We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the data column vector, and consider three choices of the square filter matrix $\mathbf{Q}$} (see [27, Eq. (1)]). The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [28]. The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$. The NP minimum-variance (NPMV) statistic adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions.
Normally, the NP statistic would be preferred. However, because its filter matrix $\mathbf{Q}$} contains both diagonal (autocorrelation) and off-diagonal (cross-correlation) terms, it is not robust against errors in modeling pulsar noise. In contrast, the OS and NPMV statistics use filter matrices that contain only off-diagonal terms, making them more robust. Details may be found in [27].
Figure 3: Histograms of NPMV detection statistic $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples. The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$. The upper horizontal axis is one-sided-Gaussian equivalent significance.
The values of these statistics and their false-alarm probabilities $p$ depend on the observational data and on the (inferred) pulsar-noise models under the null (CURN) hypothesis. While the data are fixed, the models are uncertain and vary across posterior samples, so we adopt a noise-marginalized approach [28, 29]. For each posterior-sample CURN universe, we compute the three detection statistics and evaluate their false-alarm probabilities analytically using the appropriate generalized $\chi^{2}$ distribution [33, 32]. We then form a histogram of the resulting $p$ values.
We first discuss the NPMV statistic, whose posterior $p$-value histogram is shown in Fig. 3. The mean and median false-alarm probabilities are $p=3.4\times 10^{-3}$ and $p=7.2\times 10^{-4}$ for free $\gamma_{gw}$ and $p=2.9\times 10^{-3}$ and $p=9.0\times 10^{-4}$ for fixed $\gamma_{gw}=13/3$. Respectively, these correspond to $2.7\sigma$, $3.2\sigma$, $2.8\sigma$, and $3.1\sigma$ equivalent Gaussian significances. The most likely range is shown in Fig. 3 as the band that contains 68% of the posterior points about the median. These correspond to significances in the range from $2.5-3.8\sigma$ for free $\gamma_{gw}$ and $2.6-3.7\sigma$ for $\gamma_{gw}=13/3$.
 statistic  significance mean median 68% range 95% range
OS $\sigma-$units: 4.0 4.5 (4.0, 5.1) (3.4, 5.6)
$\log_{10}p$: 4.4 5.5 (4.4, 6.7) (3.5, 8.0)
NP $\sigma-$units: 2.6 3.1 (2.5, 3.7) (1.9, 4.2)
$\log_{10}p$: 2.4 3.0 (2.2, 3.9) (1.5, 4.9)
NPMV $\sigma-$units: 2.7 3.2 (2.5, 3.8) (2.0, 4.5)
$\log_{10}p$: 2.5 3.1 (2.3, 4.2) (1.6, 5.4)
Table 3: Posterior distributions of the false-alarm probability $p$ for $\gamma_{gw}$ free. OS is the traditional “optimal” statistic, NP is the Neyman–Pearson statistic, and NPMV is the robust NP statistic. Ranges are centered about the median, and $p$-values are also translated into “equivalent-$\sigma$” units.
The corresponding histograms for the NP and OS false-alarm probabilities $p$ have a similar appearance, so are not shown. Instead, we summarize the most important information for all three statistics in Table 3. None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ threshold for confident detection [12]. However, it is notable that our analysis yields a traditional OS detection significance of $\approx 4.0\sigma$, which is substantially higher than the $\approx 3.2\sigma$ significance reported by NG [34]. This indicates that the “direct data combination” method is working as would be expected in the presence of an SGWB.
Another way to quantify detection confidence is through the Bayes factor (BF), defined as the ratio of the evidence for the HD model to that for the CURN model [12]. The evidence is the likelihood marginalized over the prior [57]. In our case, evaluating this marginalization integral is challenging because the parameter space is high-dimensional: $4\times 121+2=486$ dimensions when $\gamma_{gw}$ is free (or $485$ when $\gamma_{gw}$ is fixed to $13/3$). To compute the evidence, we use the generalized stepping-stone sampling method [58, 59, 60], which progressively “broadens” the sharply peaked likelihood using $K=8$ chains with K1=7K-1=7 steps, while also restricting the integration region in parameter space.
This yields $\ln(\mathrm{BF})=10.5\pm 1.0$, corresponding to $\mathrm{BF}\approx 40{\kern 0.83298pt}000$, which is substantially larger than the $\mathrm{BF}\approx 200$ previously reported by NG [9]. For a one-dimensional Gaussian random variable, this BF corresponds to a significance of $\sqrt{2\ln(\mathrm{BF})}\,\sigma\approx 4.6\sigma$ [29, Eq. (8)], which is below that required for a confident detection.
Figure 4: Reconstruction of the mean correlation between pulsar pairs as a function of angular separation, compared to the Hellings and Downs (HD) prediction (2). The vertical bars indicate the uncertainties in the estimates of the means.
Another detection confidence test is to estimate the inter-pulsar correlation pattern from the data [12]. The mean correlation $\Gamma_{ab}$ between pulsars $a$ and $b$ arising from an SGWB should follow the HD prediction [25, 26, 2]
$\Gamma_{ab}\equiv\frac{1}{2}(1+\delta_{ab})+\frac{3}{4}(1-\cos\theta)\!\left[\ln\Bigl(\frac{1-\cos\theta}{2}\Bigr)-\frac{1}{6}\right]\!.$ (2)
Here, $\theta$ is the angle between unit vectors $\hat{n}_{a}$ and $\hat{n}_{b}$ pointing from Earth to pulsars $a$ and $b$, so $\cos\theta=\hat{n}_{a}\cdot\hat{n}_{b}$. This pattern is shown in Fig. 4 as a function of $\theta$. [The $\delta_{ab}$ in (2) doubles the correlation for two pulsars with spatial separation smaller than the GW wavelength [2]. It plays no role in the Fig. 4 comparison: the 15 angular bins contain only cross-pulsar pairs $a\neq b$, for which the Kronecker-$\delta$ vanishes.]
An optimal method for “reconstructing” the HD correlation from the data has been derived [61] but is not yet fully implemented [44]. So, our reconstruction is based on the same method [54] previously used by EPTA [7] and NG [9], as implemented in [62]. Fig. 4 shows the result for the $20{\kern 0.83298pt}480$ free $\gamma_{gw}$ CURN posterior samples.
Our HD reconstruction adopts the NG [9] choice of 15 angular bins. The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin contains 484 pairs. For each posterior sample, the correlations of the pulsar pairs are weighted as in [62], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ for each bin. For each bin, we then construct and average $20{\kern 0.83298pt}480$ different Gaussian probability distributions. The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked. Fig. 4 shows their means (very close to medians) and standard deviations. The reconstruction matches the HD curve: the reduced $\chi^{2}$ of $0.84$ is consistent with normal statistical fluctuations ($p$-value $0.63$ for 15 degrees of freedom).
Conclusion — We have conducted a search for a stochastic gravitational-wave background (SGWB) using data from five PTAs. The analysis is straightforward in the PTA context: we apply only minor adjustments to the pulsar noise models and discard a small fraction of data. In essence, we repeat an established analysis using a larger dataset.
Our results provide the strongest evidence to date that an SGWB contributes to the observed pulsar pulse arrival-time fluctuations. However, the statistical significance of this evidence—assessed using multiple frequentist and Bayesian methods—remains too low to support a definitive detection.
Our use of existing data-analysis pipelines and infrastructures makes it straightforward to compare our results with those of the individual PTAs, but it has drawbacks. For example, these pipelines assume that, for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent. While this is not strictly correct [61, see Conclusion], the resulting effects may be negligible or readily corrected [43]. Similarly, our reconstruction of the HD curve is suboptimal because the implementation of an optimal method [61] is still in progress [44].
As part of its upcoming data release, the IPTA is re-analyzing the five-PTA plus CPTA dataset to produce an IPTA-DR3 release (new .par and .tim files). If the primary outcome of this re-analysis is the adoption of common astrophysical models for the pulsars, then—per the arguments above—we expect an SGWB search of the five-PTA data to yield results consistent with ours. But if the re-analysis reveals that some data are incorrect and must be discarded, the results could differ. In this sense, the reliability of our conclusions depends on the correctness of the five existing PTA analyses.
We have deliberately kept this analysis as straightforward as possible and have not “iterated” it. This departs from standard practice in pulsar astronomy, where models—for example, pulsar timing parameters—are iteratively refit for improvement. Here, we avoid such iterative refinement and tuning of the full PTA search pipeline. Past experience with weak-signal detection in the LIGO context, as well as theoretical studies in the PTA context [11], suggests that such iteration can be misleading.
Because they are not public, we have not been able to include CPTA data [8] in our analysis. The CPTA data come from the Five-hundred-meter Aperture Spherical Telescope (FAST) in Guizhou Province, China. FAST is the world’s most sensitive radio telescope, and its data are of very high quality. Including them in future analyses would be both interesting and informative.
Note: the complete code used for this work will be posted publicly at the time of publication.
Acknowledgments – We thank Rutger van Haasteren for many useful conversations and discussions, for his encouragement with this project, and for his comments on various drafts of the manuscript. We also thank Stanislav Babak, Siyuan Chen, Colin Clark, Oliver Dodge, Boris Goncharov, Michael Kramer, Joseph Romano, and Michele Vallisneri for helpful feedback.

### v2

@@
 Abstract
-Using public data from five pulsar timing arrays (PTAs), we search for a stationary, isotropic, and unpolarized nHz stochastic gravitational-wave background (SGWB).
-We use pulse-time-of-arrival data from 121 pulsars, which is more informative than previous searches, carried out separately by the individual PTA collaborations using only their own data.
-For pulsars observed by more than one PTA, we employ a new “direct combination” method to merge their astrophysical models and data.
-This avoids the challenge of reconciling the different PTA timing models to obtain a single “best” model.
-A central result of our analysis is posterior probability distributions for the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency.
-While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance—assessed via a Bayesian odds ratio and through noise-marginalized false-alarm probabilities ($p$-values) for three different detection statistics—remains below the conventional $5\sigma$ threshold for a confident detection.
-We also reconstruct the inter-pulsar timing-residual correlation, as a function of the angle $\theta$ between the pulsar lines of sight.
-This is consistent with the curve predicted by Hellings and Downs (HD).
-Introduction – Pulsar timing arrays (PTAs) aim to detect low frequency gravitational waves (GWs), by searching for the effects that such waves induce on the arrival times of pulsar pulses.
-These GWs create “timing residuals” which are correlated among different pulsars [1, 2].
+Using public pulse time-of-arrival data from five pulsar timing arrays (PTAs), we search for a stationary, isotropic, and unpolarized nHz stochastic gravitational-wave background (SGWB).
+This analysis is more sensitive than previous individual PTA searches because the combined 121-pulsar dataset is about four times larger than any single PTA’s.
+For pulsars observed by multiple PTAs, we employ a new “direct combination” method to merge their astrophysical models and data.
+This avoids the challenge of reconciling different PTA timing models to obtain a single “best” model.
+A central result of our analysis is the posterior distribution of the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency.
+While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance—assessed via a Bayesian odds ratio and noise-marginalized false-alarm probabilities ($p$-values) for three detection statistics—remains below the conventional $5\sigma$ threshold for a confident detection.
+The inter-pulsar timing-residual correlation, reconstructed as a function of angle $\theta$ between the pulsar lines of sight, matches the Hellings and Downs (HD) prediction.
+Introduction – Pulsar timing arrays (PTAs) aim to detect low-frequency gravitational waves (GWs) by measuring the effects these waves induce on the arrival times of pulsar pulses.
+These GWs create “timing residuals” which are correlated between pulsars [1, 2].
 Decades-long datasets with submicrosecond timing precision yield GW amplitude (strain) sensitivities of order $A_{gw}\sim 10^{-15}$ at nHz frequencies.
-Orbiting pairs of supermassive black holes (SMBHs) are one possible source of GWs.
-If the Universe contains many such pairs, then the incoherent superposition of their GW emissions would create a stochastic GW background (SGWB) [3, 4, 5] with a power-law spectrum in the nHz band.
+Supermassive black hole binaries (SMBHBs) are one source of GWs.
+If the Universe contains many of them, the incoherent superposition of their GW emissions would create a stochastic GW background (SGWB) [3, 4, 5] with a power-law spectrum in the nHz band.
 Different PTA collaborations have reported varying levels of evidence for a nHz SGWB.
-The Parkes PTA (PPTA [6]) observed 30 pulsars over 18 years, concluding that the data showed “no support for or against”.
-The European and Indian PTAs (EPTA and InPTA [7]) observed 25 pulsars, and reported “marginal evidence” in the 25-year dataset and “evidence” when only the last 10 years of data were analyzed.
-The Chinese PTA (CPTA [8]) observed 57 pulsars over three years, and reported “some evidence”.
-The North American Nanohertz Observatory for Gravitational Waves (NG [9]) observed 67 pulsars over 15 years and reported “compelling evidence”.
-Most recently, the MeerKAT PTA (MPTA [10]) observed 83 pulsars over 4.5 years, and, depending on assumptions about pulsar noise processes, reported either “high significance” or “no significance” (also see [11]).
-None of these pass the threshold of $5\sigma$ significance required to claim detection [12].
-Figure 1: A Mollweide projection showing sky locations of the 121 pulsars included in our analysis. Colors indicate how many PTAs contributed data for each pulsar.
-All of these PTAs (apart from CPTA) have released their data to the public.
-Using this full public dataset increases the total number of distinct pulsars to $N=121$, as shown in Fig.
+The Parkes PTA (PPTA [6]) observed 30 pulsars over 18 yr, concluding that the data show “no support for or against”.
+The European and Indian PTAs (EPTA and InPTA [7]) observed 25 pulsars and reported “marginal evidence” in the 25-yr dataset and “evidence” when only the last 10 yr were analyzed.
+The Chinese PTA (CPTA [8]) observed 57 pulsars over 3 yr, and reported “some evidence”.
+The North American Nanohertz Observatory for Gravitational Waves (NG [9]) observed 67 pulsars over 15 yr and reported “compelling evidence”.
+Most recently, the MeerKAT PTA (MPTA [10]) observed 83 pulsars over 4.5 yr, and, depending on assumptions about pulsar noise processes, reported either “high significance” or “no significance” (see also [11]).
+None reach the conventional $5\sigma$ significance threshold required to claim a detection [12].
+Figure 1: A Mollweide projection showing sky locations of the 121 pulsars included in our analysis. Colors indicate how many PTAs provided data for that pulsar.
+All of these PTAs (apart from CPTA) have released their data.
+Using this full public dataset increases the total number of distinct pulsars to $N_{p}=121$ as shown in Fig.
 1.
-PTAs search for an SGWB by correlating pairs of pulsars, whose number scales as $N^{2}$.
-Furthermore, since most pulsars are observed by more than one collaboration, there are also more data samples per pulsar.
-Hence, analysis of the full dataset should significantly increase the GW strain sensitivity and the detection confidence [13, 14, 15, 2].
-We analyze this five-PTA dataset [16, 17, 18, 19, 20] using a recently proposed “direct combination” method [21, 22].
-This allows us to combine the data with only minor modifications, as described below.
-Tests of this method, carried out with the International Pulsar Timing Array (IPTA) Data Release 2 (DR2) combined data set from 2019 [14], suggest that it produces results that are identical to traditional (but more time-consuming) data combination techniques.
+PTAs search for an SGWB by correlating pairs of pulsars, whose number scales as $N_{p}^{2}$.
+Furthermore, since many pulsars are observed by more than one collaboration (on average, 1.8) there are also more data samples per pulsar.
+Hence, analysis of the combined dataset should significantly increase the GW strain sensitivity and the detection confidence [13, 14, 15, 2].
+We analyze this five PTA dataset [16, 17, 18, 19, 20] using a recently proposed “direct combination” method [21, 22].
+This allows us to combine the data with minor modifications, as described below.
+Tests of this method [21], carried out with the International Pulsar Timing Array [23] (IPTA) Data Release 2 (DR2) combined dataset from 2019 [14], suggest that it produces results identical to traditional (but more time-consuming) data combination techniques.
 Aside from this, our analysis is standard.
 We use the same tools and methods as the individual PTAs, as implemented in Enterprise [24] and in the next-generation PTA data-analysis package Discovery [25].
-Both are based on parameterized likelihood models and employ Markov chain Monte Carlo methods.
-These produce posterior samples (here called “posteriors”) that approximate the inferred posterior probability distributions of the model parameters.
+Both are based on parameterized likelihood models, and employ Markov chain Monte Carlo (MCMC) methods to produce posterior samples (here called “posteriors”) which approximate the inferred posterior probability distributions of the model parameters.
+Detection confidence is assessed using both Bayesian and frequentist methods.
+Note: additional tables and plots may be found in the Supplementary Material (SM).
 Methods, Models, and Conventions – Because we use standard methods, this paper is brief.
-When we follow the PTA analyses described above [6, 7, 8, 9, 10], we give only a brief description and cite the relevant literature.
+When we follow the PTA analyses described above [6, 7, 8, 9, 10], we give only a cursory description and cite the relevant literature.
 If we do something different, then more details are provided.
@@
 Since our null hypothesis is defined via our signal hypothesis, it has exactly the same parameters/noise-models as the signal hypothesis.
-The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the HD correlations [27].
+The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the inter-pulsar HD correlations [28], hence the name “uncorrelated”.
 It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms.
@@
 Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2].
-CURN is considered a more conservative null hypothesis than that of “no GWs”.
+CURN is a more “conservative” null hypothesis than that of “no GWs”.
 Using Discovery [25] as a tool, and conditioning the Gaussian likelihood on the data and the priors, we generate posteriors for the HD and CURN models.
-These are sets of parameter values that are most the consistent with the data and priors; they characterize the joint posterior probability distributions of the model parameters.
-For this work, the most important model parameters are the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the SGWB power-law power spectrum.
-To define these, let ${\rm{d}}\rho_{gw}$ denote the SGWB energy density in the frequency range $f,f+{\rm{d}}f]$ and let $\rho_{cr}=3c^{2}H_{0}^{2}/8\pi G\simeq 7.7\times 10^{-10}\ \mathrm{J}/\mathrm{m}^{3}$ denote the critical (closure) energy density of the Universe.
+These are sets of inferred parameter values most consistent with the data and priors, thereby characterizing the posterior distributions of the model parameters.
+The most important model parameters are $A_{gw}$ and $\gamma_{gw}$, which control the power-law amplitude and exponent for the SGWB power spectrum.
+To define these, let ${\rm{d}}\rho_{gw}$ denote the SGWB energy density in the frequency range $f,f+{\rm{d}}f]$ and $\rho_{cr}=3c^{2}H_{0}^{2}/8\pi G\simeq 7.7\times 10^{-10}\ \mathrm{J}/\mathrm{m}^{3}$ denote the critical (closure) energy density of the Universe.
 Here, $c$ is the speed of light, $G$ is Newton’s gravitational constant, and $H_{0}\simeq 67.4~\mathrm{km}/\mathrm{s\,Mpc}\simeq 2.2\times 10^{-18}\,\mathrm{Hz}$ is the current Hubble expansion rate.
@@
 where $f_{r}=1/\mathrm{yr}\simeq 31.7\,\mathrm{nHz}$ is a standard reference frequency.
-Together with our previous assumptions, this spectral model (assumed to apply in the range $2<f/\mathrm{nHz}<200$) fully characterizes the SGWB.
+Together with our previous assumptions, this spectral model fully characterizes the SGWB over the 14 frequency bins of our analysis $2\lesssim f/\mathrm{nHz}\lesssim 30$.
 One result of this analysis is posterior probability distributions for $A_{gw}$ and $\gamma_{gw}$.
 We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to take the value $13/3$.
-The latter would be expected from a population of SMBHs in circular binaries, provided that GWs dominate the energy loss at nHz frequencies.
+The latter would be expected from a population of SMBHBs in circular orbits, provided that GWs dominate the energy loss at nHz frequencies.
 One way to assess detection confidence is via frequentist detection statistics [28].
@@
 All three statistics $D$ are quadratic functions of the timing residuals, with weights determined by $\bf C$ and $\bf C_{0}$.
-Each of the posteriors corresponds to a different model universe, with its own amplitudes and exponents for the different noise sources, and thus with its own covariance matrices $\bf C$ and $\bf C_{0}$.
-This means that the values of the detection statistics $D$, and the probability ($p$-value) of obtaining them under CURN, differ from one posterior sample to the next.
-Thus, to assess detection confidence, we present the posterior distributions of these $p$-values.
-In the literature, its mean is called the “noise-marginalized false-alarm probability” [28] or “posterior predictive $p$-value” [29].
-Some past work uses ad-hoc methods for assessing the false-alarm probability $p$.
+Each posterior corresponds to a different model universe, with its own amplitudes/exponents for the different noise sources, and thus its own covariance matrices $\bf C$ and $\bf C_{0}$.
+Hence, the values of the detection statistics $D$, and the probability ($p$-value) of obtaining them under CURN, differ from one posterior sample to the next.
+To assess detection confidence, we present the posterior distributions of these $p$-values.
+In the literature, its mean is called the “posterior predictive $p$-value” [29], similar in spirit to the “noise-marginalized optimal statistic” [30].
+Some past work uses ad hoc methods for assessing the false-alarm probability $p$.
 “Sky scrambling” and “phase shifting” were intended to “remove any signal from the data”.
@@
 For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
-We determine the false-alarm probability $p$ using a correct method [33, 34, 32, 27].
-Since $D$ is a quadratic form in the data, its distribution under CURN is a generalized $\chi^{2}$ distribution [33, 32, 27]; the probability that it exceeds the observed value in one of our posterior samples is found analytically.
+We determine the false-alarm probability $p$ using a statistically rigorous method [34, 35, 33, 28].
+Since $D$ is a quadratic form in the data, its distribution under CURN is a generalized $\chi^{2}$ distribution [34, 33, 28]; the probability that it exceeds the observed value in each of our posterior samples is found analytically.
 To aid interpretation, we also report the $p$-values as “equivalent $q\sigma$” levels: $q>0$ is chosen so that a random draw from a zero-mean Gaussian with variance $\sigma^{2}$ has probability $p$ of exceeding $q\sigma$.
@@
 The latter, a false-alarm probability of less than one per three million, is the traditional threshold for confident detection [12].
-Analysis of the five-PTA dataset – The data from the five PTAs come from 121 millisecond pulsars, as shown in Fig.
+Analysis of the five PTA dataset – The data from the five PTAs come from 121 millisecond pulsars, as shown in Fig.
 1.
-These consist of the second EPTA data release (10-year subset) [35, 16], the first InPTA data release [36, 17], the MPTA 4.5-yr data release [37, 18], the NG 15yr data release [38, 19], and the third PPTA data release [39, 20].
-Each PTA’s data consists of one .par file and one or more .tim files, per pulsar.
+These consist of the second EPTA data release (10-yr subset) [36, 16], the first InPTA data release [37, 17], the MPTA 4.5-yr data release [38, 18], the NG 15-yr data release [39, 19], and the third PPTA data release [40, 20].
+Each PTA’s public open dataset has one .par file and one or more .tim files per pulsar.
+The data we analyze consist of $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ times of arrival (TOAs): the complete set of observations from 25 EPTA pulsars, 14 InPTA pulsars, 83 MPTA pulsars, 68 NG pulsars, and 32 PPTA pulsars.
 We merge these using the direct combination method [21, 22].
-The .tim files, containing pulse times-of-arrival (TOAs) and their uncertainties, are used “as is”.
-The .par files, containing the astrophysical models, and some detector-specific parameters, are modified (in effect) as follows, one pulsar at a time.
-(This is also the first step of “manual” data combination.)
+If not already present, rotation counts are added to the .tim files, which contain pulse TOAs and uncertainties.
+The .par files, containing astrophysical models and detector-specific parameters, are modified one pulsar at a time (the first step of “manual” data combination), as follows.
   1. 1.
     For pulsars observed by more than one PTA, a “reference PTA” is selected via the arbitrary ordering $\rm{NG}>\rm{EPTA}>\rm{PPTA}>\rm{MPTA}$.
-(InPTA is not listed because its pulsars are observed by other PTAs.) Any other PTAs that have data for this pulsar are called “target PTAs”.
+InPTA is not listed because its pulsars are observed by other PTAs.
+Other PTAs that have data for this pulsar are called “target PTAs”.
   2. 2.
@@
   3. 3.
-    Any .par file that uses the TCB time standard is converted to TDB using Tempo2 [40] tools.
+    If the pulsar’s .par files use both TCB and TDB time conventions, then quantities whose units contain time are converted to TDB with Tempo2 [41] tools. See SM Sec. III.
   4. 4.
-    If the reference .par file has a DMX model for DM fluctuations (NG and InPTA only), then it is replaced by a Gaussian process (DMGP) model.
-The parameter DM0 is initialized to the mean DMX value, and DM1 and DM2 are initialized to zero.
+    The reference .par file DM model is replaced with a constant-in-time DM model: DM0 is initialized to the .par file DM value at time DMEPOCH, and DM1 and DM2 are initialized to zero.
   5. 5.
-    If the reference .par file has a DMGP model, then the DM0 value is retained, and DM1 and DM2 are initialized to zero.
-  6. 6.
-    The astrophysical model $\beta$ [1, Eq.
-7.2], including sky position, proper motion, spin frequency, spin-down parameters, binary orbital parameters, and DM values, is copied from the reference .par file to the target .par files.
-Detector-specific parameters in the .par files, for example, “JUMP” for time offsets and “FDx” for narrow-band frequency offsets, are unchanged.
-This allows for template model shape differences between PTAs.
-For four pulsars, ad-hoc “tricks” are used to maintain phase connection between the .tim file TOAs.
-For the binary pulsar PSR J2241-5236 (MPTA, PPTA), we retain the MPTA 11-parameter model and the PPTA 18-parameter model for the orbital frequency evolution.
-The other parameters follow 6 above.
-For PSR J0610-2100 (MPTA, NG) and PSR J1024-0719 (EPTA, MPTA, NG, PPTA), we keep the DMX model for the NG data and the DMGP model for the remaining PTAs.
-For PSR J1804-2858 (MPTA), we retain the values of DM1 and DM2 rather than initializing them to zero.
-The direct combination method [21, 22] works because the only significant change it makes is to the parameters $\beta$, item 6 above.
-While their exact values are unknown, every pulsar has unique values for these parameters, for example, the sky position.
-Those enter the analysis via a nominal value $\beta_{0}$ [1, Eq.
-7.2] and via a matrix $\mathbf{M}$} [1, Eq.
-7.4], which is the rate of change of the pulse arrival-time with respect to variations $\epsilon$ in $\beta$.
-In the (tiny) region of interest, $\mathbf{M}$} varies slowly with those parameters, so computing $\mathbf{M}$} with the same $\beta_{0}$ for a given pulsar in all PTAs introduces negligible errors.
+    A nominal astrophysical model $\beta_{0}$ [1, Eq.
+7.2] (sky position, proper motion, spin frequency, spin-down parameters, binary orbital parameters, and DM values) is copied from the reference .par file to the target .par files.
+Detector-specific parameters in the .par files, for example “FDx” for narrow-band frequency offsets, are unchanged.
+One additional adjustable phase shift parameter (“JUMP”) per target PTA is introduced to account for pulse profile template shape differences.
+The direct combination method [21, 22] adopts one set of parameter values $\beta_{0}$ for each pulsar, item 5 above.
+It works because each pulsar has exact but unknown values $\beta$ for these (for example, for the sky position).
+They enter the analysis via the nominal value $\beta_{0}$ [1, Eq.
+7.2] and via a “design matrix” $\mathbf{M}$} [1, Eq.
+7.4], which is the rate of change of the pulse arrival-time with respect to variations in $\beta$, evaluated at $\beta_{0}$.
+In the (tiny) region of interest, $\mathbf{M}$} varies slowly with those parameters, so computing $\mathbf{M}$} at the same nominal $\beta_{0}$ for a given pulsar introduces negligible errors.
 The subsequent marginalization over $\epsilon=\beta-\beta_{0}$ [1, Eq.
 7.6] fully accounts for the uncertainties in $\beta$ and for the variations in its nominal value across PTAs.
-Similarly, the RN and DM noise are intrinsic to the pulsar and to the interstellar medium (ISM), so must be the same for all PTAs.
+Similarly, the RN and DM noise are intrinsic to the pulsar and to the interstellar medium (ISM), so must be the same across PTAs.
+We therefore adopt the physically motivated DMGP model [42], used by all PTAs except NG and InPTA, which use the DMX model.
+The design matrix $\mathbf{M}$} allows the effects of DM fluctuations on the timing residuals to be computed from the fixed-in-time DM timing residuals.
 The next analysis step is single-pulsar noise modeling.
-The goal is to determine each pulsar’s WN covariance matrix $\mathbf{N}$}, which appears in the Gaussian likelihood [1, Eq.
+These determine each pulsar’s WN covariance matrix $\mathbf{N}$}, which appears in the Gaussian likelihood [1, Eq.
 7.27].
-This characterizes the timing residuals fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eq.
-7.2,7.26].
+This characterizes the timing residual fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eqs.
+7.2 and 7.26], modeled with 30 and 100 frequency bins respectively.
 Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
@@
 The WN covariance matrix $\mathbf{N}$} is set by these parameters and by the .tim file contents.
-NG, PPTA, MPTA, and seven EPTA backends use the parameters $EFAC$, EQUADEQUAD, and $ECORR$ [38, Eq.
+NG, PPTA, MPTA, and seven EPTA backends use the parameters $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and $\mathrm{ECORR}$} [39, Eq.
 2].
-InPTA and the remaining EPTA backends only use $EFAC$ and $EQUAD$ [41].
+InPTA and the remaining EPTA backends only use $\mathrm{EFAC}$} and $\mathrm{EQUAD}$} [43].
 Parameters Uniform on range Stage
@@
 $\log_{10}A$ (-20, -11) S, F
-Table 1: Priors for the WN parameters, and for the amplitudes/exponents of the power-law RN, DM and SGWB noise spectra. “S” denotes single pulsar, and “F” denotes full PTA.
-The WN parameters for a given pulsar are obtained from all of its TOA data (meaning, all .tim files).
-These are analyzed using JAXopt [42], with priors taken from the Table 1 rows with an “S” in the final column; the amplitude and exponent priors are used for all power laws.
-For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model; for $\gamma>1$, the $f\to\infty$ divergences are absent [43, 44].
-Note that $EQUAD$ has the units used in Tempo [45], Tempo2 [40] and PINT [46], and is often written as $T2EQUAD$.
-The output is a “WN dictionary” of best-fit $EFAC$, EQUADEQUAD and (potentially) $ECORR$ values, which are used in subsequent analysis to obtain $\mathbf{N}$} [38, Eq.
+Table 1: Priors for the WN parameters, and for the amplitudes/exponents of the power-law RN, DM and SGWB noise spectra.
+“S” denotes single pulsar analysis; “F” denotes multipulsar analysis.
+The WN dictionary found in the “S” stage is used in the “F” stage.
+The WN parameters for a given pulsar are obtained from all of its (.tim files) TOA data.
+These are analyzed using JAXopt [44], with priors taken from the Table 1 rows marked “S”; the amplitude/exponent priors are used for all power laws.
+For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model, while for $\gamma>1$ the $f\to\infty$ divergences are absent [45, 46].
+Note that $\mathrm{EQUAD}$} has the units used in Tempo [47], Tempo2 [41], and PINT [48], and is written as $\rm T2EQUAD$.
+The output is a “WN dictionary” of best-fit $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and (potentially) $\mathrm{ECORR}$} values, used later to obtain $\mathbf{N}$} [39, Eq.
 2].
-For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA-analysis values.
+For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA analysis values.
 This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more reliable estimates of the WN parameters.
-We drop TOAs for any backends which have $EFAC<0.1$ or $EFAC>10$, since this indicates that the backend or template matching was unstable, or that RFI or calibration or instrumental issues were present.
-This removes $0.17\%$ of the TOAs.
-Our final data for analysis comes from 25 EPTA pulsars, 13 (of 14) InPTA pulsars, 83 MPTA pulsars, 68 NG pulsars, and 31 (of 32) PPTA pulsars.
-(Data from InPTA pulsar J1022+1001 are dropped because all the $EFAC$s are larger than 10.
-Data from PPTA pulsar PSR J1741+1351 are dropped because there were only 16 observations.)
+In SM Sec.
+I, we show that our conclusions are robust: they are not changed much by dropping TOAs from backends with the largest $\mathrm{EFAC}$} corrections.
 The final stage is multi-pulsar analysis, following the procedures described in [1, 9].
@@
 The use of Discovery has significantly sped up our analysis.
-Within the Discovery framework, we use the priors listed in Tab.
+Within the Discovery framework, we use the WN dictionary plus the SGWB priors listed in Tab.
 1 with an “F” in the final column.
@@
 NUTS is the default for Discovery [25], which provides the log-likelihood gradients needed by Hamiltonian samplers.
-It is much faster than the PTMCMC sampler [53] used by the individual PTAs when they analyzed their own data.
-We generated $20{\kern 0.83298pt}480$ posterior samples under the HD signal hypothesis, and the same number of posteriors under the CURN hypothesis.
+It generates samples that are much more independent than the parallel tempering MCMC sampler [55] previously used by the individual PTAs, and is also much faster.
+We generated $20{\kern 0.83298pt}480$ posterior samples under the HD signal hypothesis, and the same number under the CURN hypothesis.
 Each gives the amplitudes and exponents of the power-law models describing the pulsar (red and DM) noise, and the SGWB.
@@
 Figure 2: The posteriors for the amplitude and exponent of the SGWB power spectrum, as defined in (1).
-Results of the five-PTA analysis – One important result of this analysis are the posterior distributions of the SGWB power-spectrum amplitude $A_{gw}$ and exponent $\gamma_{gw}$ as defined in (1).
+Results of the five PTA analysis – One important result of this analysis are the posterior distributions of the SGWB power-spectrum amplitude $A_{gw}$ and exponent $\gamma_{gw}$ as defined in (1).
 These are shown in Fig.
@@
 As expected, the posteriors are very similar for the two hypotheses.
-This is because, if there are $\lesssim 1500$ pulsars (see [54, Fig.
-10]), most of the SGWB evidence arises from autocorrelations rather than cross-correlations—–and by construction, CURN includes these autocorrelations.
+This is because, if there are $\lesssim 1500$ pulsars (see [56, Fig.
+10]), most of the SGWB information is provided by auto- rather than cross-correlations, and by construction, CURN includes these autocorrelations.
 Model $\log_{10}A_{gw}$ $\gamma_{gw}$
-CURN with free $\gamma_{gw}\qquad$ $14.24^{+0.11}_{-0.11}\qquad$ $3.30^{+0.28}_{-0.27}$
+CURN with free $\gamma_{gw}\qquad$ $14.37^{+0.11}_{-0.12}\qquad$ $3.63^{+0.27}_{-0.26}$
 CURN with $\gamma_{gw}=13/3\qquad$ $14.69^{+0.04}_{-0.04}\qquad$
-HD with free $\gamma_{gw}\qquad$ $14.28^{+0.12}_{-0.12}\qquad$ $3.32^{+0.28}_{-0.28}$
-HD with $\gamma_{gw}=13/3\qquad$ $14.72^{+0.05}_{-0.05}\qquad$
-Table 2: Median posteriors values for $A_{gw}$ and $\gamma_{gw}$, and their 68% containment intervals (gray bands in Fig. 3).
-Happily, these posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6].
+HD with free $\gamma_{gw}\qquad$ $14.45^{+0.13}_{-0.15}\qquad$ $3.76^{+0.31}_{-0.29}$
+HD with $\gamma_{gw}=13/3\qquad$ $14.72^{+0.04}_{-0.04}\qquad$
+Table 2: Median posterior values for $A_{gw}$ and $\gamma_{gw}$, and their 68% containment intervals (gray bands in Fig. 3).
+These posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6].
 They are more constraining than the posteriors obtained by the individual PTAs: the uncertainty ellipses in the bottom-left plot of Fig.
-2 are smaller than those obtained by intersecting the individual PTA posteriors [55].
-For example, the $\text{area}=\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse (for the HD hypothesis) is 45% of that shown in [55, Fig.
+2 are smaller (by about a factor of two) than those obtained by intersecting the individual PTA posteriors [57].
+For example, the $\text{area}\equiv\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse for the HD hypothesis is 48% of that shown in [57, Fig.
 1, right].
 As explained earlier, this is as expected for an SGWB: joint analysis yields more evidence and is more constraining than naive intersection of the separate PTA results.
-To claim an SGWB detection, there are several hurdles to overcome [12].
-One must show that the data is consistent with the HD signal hypothesis, and that it is very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis.
-Clearly, the estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are almost identical under both hypotheses.
-One way to assess significance is through null-hypothesis testing [56] using a frequentist statistic.
-We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the data column vector, and consider three choices of the square filter matrix $\mathbf{Q}$} (see [27, Eq.
+The PTA community has established basic criteria for claiming a detection [12].
+Among these, one must show that the data is consistent with the HD signal hypothesis, and very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis.
+The estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are very similar under both hypotheses.
+One way to assess significance is the frequentist approach: null-hypothesis testing [58] using a statistic.
+This computes the (noise-marginalized false-alarm) probability $p$ that signal-like data would occur under the null hypothesis.
+We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the data column vector, and consider three choices of the square filter matrix $\mathbf{Q}$} (see [28, Eq.
 (1)]).
-The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [28].
-The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$.
-The NP minimum-variance (NPMV) statistic adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions.
+The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [30, 33, 28].
+The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$ [28].
+The NP minimum-variance (NPMV) statistic [28] adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions.
+Why?
 Normally, the NP statistic would be preferred.
@@
 The upper horizontal axis is one-sided-Gaussian equivalent significance.
-The values of these statistics and their false-alarm probabilities $p$ depend on the observational data and on the (inferred) pulsar-noise models under the null (CURN) hypothesis.
-While the data are fixed, the models are uncertain and vary across posterior samples, so we adopt a noise-marginalized approach [28, 29].
-For each posterior-sample CURN universe, we compute the three detection statistics and evaluate their false-alarm probabilities analytically using the appropriate generalized $\chi^{2}$ distribution [33, 32].
-We then form a histogram of the resulting $p$ values.
+The detection statistic values and their false-alarm probabilities $p$ depend on the observational data and (inferred) pulsar-noise models under the null (CURN) hypothesis.
+While the data are fixed, the models are uncertain and vary across posteriors, so we adopt a noise-marginalized approach [30, 29].
+For each posterior-sample CURN universe, we compute the three statistics and evaluate their false-alarm probabilities using the corresponding generalized-$\chi^{2}$ distribution [34, 33].
+We then form a histogram of the resulting $p$-values.
 We first discuss the NPMV statistic, whose posterior $p$-value histogram is shown in Fig.
 3.
-The mean and median false-alarm probabilities are $p=3.4\times 10^{-3}$ and $p=7.2\times 10^{-4}$ for free $\gamma_{gw}$ and $p=2.9\times 10^{-3}$ and $p=9.0\times 10^{-4}$ for fixed $\gamma_{gw}=13/3$.
-Respectively, these correspond to $2.7\sigma$, $3.2\sigma$, $2.8\sigma$, and $3.1\sigma$ equivalent Gaussian significances.
+The mean/median false-alarm probabilities are $p=5.5\times 10^{-4}$$p=1.1\times 10^{-4}$ for free $\gamma_{gw}$ and $p=3.7\times 10^{-4}$$p=8.0\times 10^{-5}$ for fixed $\gamma_{gw}=13/3$.
+Respectively, these correspond to $3.3\sigma$$3.7\sigma$ and $3.4\sigma$$3.8\sigma$ equivalent Gaussian significances.
 The most likely range is shown in Fig.
 3 as the band that contains 68% of the posterior points about the median.
-These correspond to significances in the range from $2.5-3.8\sigma$ for free $\gamma_{gw}$ and $2.6-3.7\sigma$ for $\gamma_{gw}=13/3$.
+These correspond to significances in the range from $3.1\sigma$ to $4.3\sigma$ for free $\gamma_{gw}$ and $3.3\sigma$ to $4.3\sigma$ for $\gamma_{gw}=13/3$.
  statistic  significance mean median 68% range 95% range
-OS $\sigma-$units: 4.0 4.5 (4.0, 5.1) (3.4, 5.6)
-$\log_{10}p$: 4.4 5.5 (4.4, 6.7) (3.5, 8.0)
-NP $\sigma-$units: 2.6 3.1 (2.5, 3.7) (1.9, 4.2)
-$\log_{10}p$: 2.4 3.0 (2.2, 3.9) (1.5, 4.9)
-NPMV $\sigma-$units: 2.7 3.2 (2.5, 3.8) (2.0, 4.5)
-$\log_{10}p$: 2.5 3.1 (2.3, 4.2) (1.6, 5.4)
+OS $\sigma-$units: 4.3 4.8 (4.3, 5.4) (3.9, 5.9)
+$\log_{10}p$: 5.2 6.2 (5.1, 7.4) (4.3, 8.7)
+NP $\sigma-$units: 3.3 3.8 (3.2, 4.3) (2.7, 4.9)
+$\log_{10}p$: 3.3 4.1 (3.2, 5.1) (2.4, 6.3)
+NPMV $\sigma-$units: 3.3 3.7 (3.1, 4.3) (2.7, 4.9)
+$\log_{10}p$: 3.3 4.0 (3.1, 5.1) (2.4, 6.3)
 Table 3: Posterior distributions of the false-alarm probability $p$ for $\gamma_{gw}$ free.
@@
 Ranges are centered about the median, and $p$-values are also translated into “equivalent-$\sigma$” units.
-The corresponding histograms for the NP and OS false-alarm probabilities $p$ have a similar appearance, so are not shown.
-Instead, we summarize the most important information for all three statistics in Table 3.
-None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ threshold for confident detection [12].
-However, it is notable that our analysis yields a traditional OS detection significance of $\approx 4.0\sigma$, which is substantially higher than the $\approx 3.2\sigma$ significance reported by NG [34].
-This indicates that the “direct data combination” method is working as would be expected in the presence of an SGWB.
-Another way to quantify detection confidence is through the Bayes factor (BF), defined as the ratio of the evidence for the HD model to that for the CURN model [12].
-The evidence is the likelihood marginalized over the prior [57].
-In our case, evaluating this marginalization integral is challenging because the parameter space is high-dimensional: $4\times 121+2=486$ dimensions when $\gamma_{gw}$ is free (or $485$ when $\gamma_{gw}$ is fixed to $13/3$).
-To compute the evidence, we use the generalized stepping-stone sampling method [58, 59, 60], which progressively “broadens” the sharply peaked likelihood using $K=8$ chains with K1=7K-1=7 steps, while also restricting the integration region in parameter space.
-This yields $\ln(\mathrm{BF})=10.5\pm 1.0$, corresponding to $\mathrm{BF}\approx 40{\kern 0.83298pt}000$, which is substantially larger than the $\mathrm{BF}\approx 200$ previously reported by NG [9].
-For a one-dimensional Gaussian random variable, this BF corresponds to a significance of $\sqrt{2\ln(\mathrm{BF})}\,\sigma\approx 4.6\sigma$ [29, Eq.
-(8)], which is below that required for a confident detection.
+The corresponding histograms for the NP and OS false-alarm probabilities $p$ are similar (see SM Sec.
+II).
+We list key results for all three statistics in Table 3.
+None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ detection threshold [12].
+However, note that our analysis yields a traditional OS detection significance of $\approx 4.3\sigma$, which is higher than the $\approx 3.2\sigma$ significance reported by NG [35].
+This suggests that the “direct data combination” method is working as would be expected in the presence of an SGWB.
+Another way to quantify detection confidence [12] is via the Bayes factor ${\mathrm{BF}}=E_{\rm{HD}}/E_{\rm{CURN}}$, the ratio of HD to CURN model evidences.
+Here, the evidence $E$ is the likelihood marginalized over the prior [59]; $E$ has units $({\rm time})^{-N}$ where $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ is the number of TOAs (see SM Sec.
+I Table 4).
+Evaluating these integrals is challenging because the parameter space has $4\times 121+2=486$ dimensions for $\gamma_{gw}$ free ($485$ for $\gamma_{gw}=13/3$).
+To compute the evidences we use generalized stepping-stone sampling [60, 61, 62], which broadens the sharply peaked likelihood using $K=8$ chains (K1=7K-1=7 steps) while restricting the integration region.
+This gives $\ln(E_{\rm{HD}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}504.39\pm 0.10$ and $\ln(E_{\rm{CURN}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}494.21\pm 0.09$, implying $\ln({\mathrm{BF}})=10.18\pm 0.13$ and ${\mathrm{BF}}=26{\kern 0.83298pt}000\pm 3{\kern 0.83298pt}000$.
+This is two orders-of-magnitude larger than the Bayes factor ${\mathrm{BF}}\approx 200$ reported by NG [9].
+For a real Gaussian random variable, the corresponding significance [29, Eq.
+(8)] of $\sqrt{2\ln({\mathrm{BF}})}\,\sigma\approx 4.5\sigma$ is below the threshold for confident detection.
 Figure 4: Reconstruction of the mean correlation between pulsar pairs as a function of angular separation, compared to the Hellings and Downs (HD) prediction (2).
 The vertical bars indicate the uncertainties in the estimates of the means.
-Another detection confidence test is to estimate the inter-pulsar correlation pattern from the data [12].
+We test if the inter-pulsar correlations match those expected from an SGWB by estimating the angular correlation pattern from the data [12].
 The mean correlation $\Gamma_{ab}$ between pulsars $a$ and $b$ arising from an SGWB should follow the HD prediction [26, 27, 2]
@@
 4 shows the result for the $20{\kern 0.83298pt}480$ free $\gamma_{gw}$ CURN posterior samples.
-Our HD reconstruction adopts the NG [9] choice of 15 angular bins.
-The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin contains 484 pairs.
-For each posterior sample, the correlations of the pulsar pairs are weighted as in [62], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ for each bin.
-For each bin, we then construct and average $20{\kern 0.83298pt}480$ different Gaussian probability distributions.
+Our HD reconstruction uses the NG [9] choice of 15 angular bins.
+The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin has 484 pairs.
+For each posterior sample, the correlations of the pulsar pairs are weighted as in [64], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ per bin.
+For each bin, we then construct and average the $20{\kern 0.83298pt}480$ corresponding Gaussian probability distributions.
 The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked.
 Fig.
-4 shows their means (very close to medians) and standard deviations.
-The reconstruction matches the HD curve: the reduced $\chi^{2}$ of $0.84$ is consistent with normal statistical fluctuations ($p$-value $0.63$ for 15 degrees of freedom).
-Conclusion — We have conducted a search for a stochastic gravitational-wave background (SGWB) using data from five PTAs.
-The analysis is straightforward in the PTA context: we apply only minor adjustments to the pulsar noise models and discard a small fraction of data.
-In essence, we repeat an established analysis using a larger dataset.
-Our results provide the strongest evidence to date that an SGWB contributes to the observed pulsar pulse arrival-time fluctuations.
-However, the statistical significance of this evidence—assessed using multiple frequentist and Bayesian methods—remains too low to support a definitive detection.
-Our use of existing data-analysis pipelines and infrastructures makes it straightforward to compare our results with those of the individual PTAs, but it has drawbacks.
-For example, these pipelines assume that, for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent.
-While this is not strictly correct [61, see Conclusion], the resulting effects may be negligible or readily corrected [43].
-Similarly, our reconstruction of the HD curve is suboptimal because the implementation of an optimal method [61] is still in progress [44].
-As part of its upcoming data release, the IPTA is re-analyzing the five-PTA plus CPTA dataset to produce an IPTA-DR3 release (new .par and .tim files).
-If the primary outcome of this re-analysis is the adoption of common astrophysical models for the pulsars, then—per the arguments above—we expect an SGWB search of the five-PTA data to yield results consistent with ours.
-But if the re-analysis reveals that some data are incorrect and must be discarded, the results could differ.
-In this sense, the reliability of our conclusions depends on the correctness of the five existing PTA analyses.
-We have deliberately kept this analysis as straightforward as possible and have not “iterated” it.
-This departs from standard practice in pulsar astronomy, where models—for example, pulsar timing parameters—are iteratively refit for improvement.
-Here, we avoid such iterative refinement and tuning of the full PTA search pipeline.
-Past experience with weak-signal detection in the LIGO context, as well as theoretical studies in the PTA context [11], suggests that such iteration can be misleading.
-Because they are not public, we have not been able to include CPTA data [8] in our analysis.
-The CPTA data come from the Five-hundred-meter Aperture Spherical Telescope (FAST) in Guizhou Province, China.
-FAST is the world’s most sensitive radio telescope, and its data are of very high quality.
-Including them in future analyses would be both interesting and informative.
-Note: the complete code used for this work will be posted publicly at the time of publication.
-Acknowledgments – We thank Rutger van Haasteren for many useful conversations and discussions, for his encouragement with this project, and for his comments on various drafts of the manuscript.
-We also thank Stanislav Babak, Siyuan Chen, Colin Clark, Oliver Dodge, Boris Goncharov, Michael Kramer, Joseph Romano, and Michele Vallisneri for helpful feedback.
+4 shows their means (close to medians) and standard deviations.
+A traditional $\chi^{2}$ goodness-of-fit test shows that the reconstruction matches the HD curve well: the reduced $\chi^{2}$ of $0.74$ is consistent with expected statistical fluctuations for 15 degrees of freedom
+Conclusion — We search for an SGWB using public data from five PTAs.
+In effect, we repeat an established analysis, but with a larger dataset.
+This is enabled by the physically motivated step of adopting consistent astrophysical parameters and noise models for all pulsars.
+Our results provide the strongest evidence to date that an SGWB contributes to observed pulsar pulse-arrival-time fluctuations.
+However, the statistical significance—–assessed using multiple frequentist and Bayesian methods—–remains too low to claim a detection.
+Using existing analysis pipelines enables straightforward comparison with individual PTA results, with which our findings are consistent, but it has limitations.
+For example, these pipelines assume that for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent.
+Although this assumption is not strictly valid [63, see Conclusion], its impact may be negligible or correctable [45].
+Likewise, our HD-curve reconstruction is suboptimal because implementation of an optimal method [63] is still in progress [46].
+Another limitation is the lack of a hierarchical pulsar-noise model [65, 66], which could shift the inferred SGWB amplitude and spectral slope [43].
+For its upcoming DR3 release, the IPTA [23] is reanalyzing the five PTA plus CPTA dataset to make new .par and .tim files.
+If the primary outcome is adopting common astrophysical models for the pulsars, then—per the arguments above—we expect an SGWB search of the DR3 five PTA data to yield results consistent with ours.
+But if reanalysis reveals that some data are incorrect and must be modified or discarded, the results could differ.
+Thus, the reliability of our conclusions depends on the correctness of the five existing PTA analyses.
+We deliberately kept this analysis straightforward and did not “iterate” it.
+This departs from standard pulsar-astronomy practice, where models (e.g., for pulsar timing) are iteratively revised.
+Here, we avoid iterative refinement and tuning of the full PTA search pipeline.
+Past experience with weak-signal detection in LIGO, as well as theoretical studies for PTAs [11], suggests that such iteration can be misleading.
+Because they are not public, we could not include CPTA data [8] in our analysis.
+These are from the Five-hundred-meter Aperture Spherical Telescope (FAST) in Guizhou, China, which is the world’s most sensitive radio telescope.
+Including these high-quality FAST data in future analyses would be interesting and informative.
+Notes: (1) Our full analysis code will be released publicly upon publication.
+(2) After this work was circulated, a different data-combination method was proposed [67].
+Unlike ours, this does not use a physically consistent timing model for a given pulsar.
+See also SM Sec.
+I.
+Acknowledgments – We thank R.
+van Haasteren for many useful conversations, encouragement on this project, and comments on manuscript drafts.
+We also thank S.
+Babak, S.
+Chen, C.
+Clark, O.
+Dodge, B.
+Goncharov, M.
+Kramer, P.M.
+Meyers, M.A.
+Papa, J.D.
+Romano, and M.
+Vallisneri for helpful feedback.
+Supplementary Material
+Here, we provide additional plots and tables as supporting material.
+I Effects discarding “outlier backend” TOAs
+In this work, we analyze the full five PTA dataset, which has $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ TOAs.
+We were curious to see if the results would change if we discarded “outlier” data.
+By this, we mean data for which the single pulsar analysis gives WN parameters differing significantly from those naively implied by the .tim file contents.
+We flag such outliers via $\mathrm{EFAC}$}, whose nominal value is (close to) unity, as shown in Fig.
+5.
+Figure 5: Histograms showing how many backends and TOAs have a given value of $\mathrm{EFAC}$}.
+To see the effects of discarding outliers, we drop TOAs from backends where $\mathrm{EFAC}$} lies outside the range $[1/Q,Q]$ for $Q=10,4,3$ and $2$.
+The number of backends/TOAs and fraction of data eliminated by these cuts are shown in Table 4.
+Table 4: Summary of EFAC cuts and their impact on backends and TOAs.
+EFAC Backends Backends TOAs TOAs % TOAs
+range discarded retained discarded retained discarded
+$0,\infty)$ 0 976 0 $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ 0
+$[0.10,10]$ 7 PPTA 969 $1{\kern 0.83298pt}915$ $1{\kern 0.83298pt}088{\kern 0.83298pt}291$ 0.18
+$[0.25,4]$ 6 InPTA, 1 NG, 19 PPTA 960 $8{\kern 0.83298pt}226$ $1{\kern 0.83298pt}081{\kern 0.83298pt}980$ 0.75
+$[0.33,3]$ 12 InPTA, 2 NG, 12 PPTA 950 $17{\kern 0.83298pt}002$ $1{\kern 0.83298pt}073{\kern 0.83298pt}204$ 1.56
+$[0.50,2]$ 1 EPTA, 24 InPTA, 4 NG, 18 PPTA 929 $21{\kern 0.83298pt}340$ $1{\kern 0.83298pt}068{\kern 0.83298pt}866$ 1.96
+The impact of dropping these “outlier” TOAs is negligible, as can be seen from the $A_{gw}$ and $\gamma_{gw}$ posteriors shown in Fig.
+6.
+These should be compared with Fig.
+2 in the paper.
+The tighter data cuts do not significantly affect the posteriors or other conclusions.
+We conclude that our “all are welcome” approach to data inclusion has not significantly skewed or affected our results, but note that discarding the “outlier” data systematically increases the inferred amplitude of the SGWB.
+Figure 6: We illustrate the effects of using different $\mathrm{EFAC}$} thresholds to eliminate data from the most poorly modeled backends.
+The black line shows the CURN posterior from Fig.
+2 in the paper.
+The other curves are constructed after dropping TOAs as shown in Table 4, and demonstrate that the CURN posteriors are not significantly affected by the data cuts.
+Note: each curve shows $20{\kern 0.83298pt}480$ posterior samples.
+A recently posted paper [67] comments about “anomalous WN parameters” in a preliminary draft of this work.
+These arose because that earlier analysis contained phase disconnections between some TOAs observed with different backends.
+This was also the reason that 5 pulsars were given “special” treatment there.
+These phase disconnection issues were fixed by adding rotation counts to the input files.
+II False-alarm probability $p$-value distributions
+In the paper, Figure 3 shows a histogram of the posterior values of the false-alarm probability for the NPMV detection statistic.
+Here, in Figures 7 and 8, we provide the corresponding posterior $p$-value histograms for the Neyman-Pearson (NP) statistic and the traditional “optimal statistic” (OS).
+The means, medians, and widths of these distributions are reported in Table 3 of the paper.
+Figure 7: Histograms of NP detection statistic $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
+The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
+The upper horizontal axis is one-sided-Gaussian equivalent significance.
+Figure 8: Histograms of the traditional optimal detection statistic (OS) $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
+The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
+The upper horizontal axis is one-sided-Gaussian equivalent significance.
+III Data Analysis Workflow
+Figure 9: A graphical illustration of the data analysis workflow.
+The input data are .par and .tim files from the five PTAs.
+From these, we produce feather files ($F$), a white noise dictionary (WND), and posterior samples (PS) which are used for further analysis.
+The circled numbers one to five order the analysis steps.
+Our data analysis workflow is illustrated in Fig.
+9.
+It begins at the top left with .par and .tim files coming from the five PTAs’ public data releases.
+The .par files contain parameters such as the pulsar sky position and spin frequency, and the .tim files contain the TOAs.
+The PTAs employ pulsar timing programs to prepare these files.
+These programs read and write the files, and are used to fit astrophysical parameters to the data, by minimizing the timing residuals.
+NG uses the PINT for this purpose, whereas the other four PTAs use Tempo2.
+The direct data combination method adopts a common astrophysical model for each pulsar.
+In practice, we do this by modifying the .par files, as described in detail in the paper.
+To analyze the data for a given pulsar, we use a single system of units.
+For example, if some of the .par files for a given pulsar have TCB time coordinates, and others have TDB time coordinates, then we convert all of them to TDB.
+Quantities that contain units of time are scaled in value by this conversion.
+For example, pulsar rotation frequencies expressed in TDB are larger than in TCB by about 16 parts per billion.
+Alternatively, if all of the .par files use TDB or TCB exclusively, then time-coordinate-dependent quantities are not modified.
+The situation is simpler for the sky position units.
+Sky positions can be specified using right ascension (RA) and declination (DEC) or in ecliptic latitude and longitude, and the angles can be given in hour-angles, degrees, or radians.
+For constructing a common timing design matrix $\mathbf{M}$}, we need to ensure that only one combination of these units is employed for a given pulsar.
+This is accomplished when the sky position is copied from the reference to the target .par files.
+For .par files that specify the ELL1H binary model, there is an additional complication.
+In PINT, if the Shapiro delay parameter $H3$ is not specified, then it defaults to zero.
+However, until a recent revision, this was incompatible with Tempo2, which did not set a default value, but instead read a random value from uninitialized memory.
+Our solution: if there is no $H3$ specified for the ELL1H model, then we set binary model to ELL1 model, which is equivalent to ELL1H with $H3=0$.
+Note that while PINT and Tempo2 internally use slightly different values for physical constants such as the solar mass, this makes no difference in our analysis.
+The modified .par files and the .tim files (with pulse counts added) are then passed to Enterprise.
+This computes the timing residuals and design matrix $\mathbf{M}$} for each pulsar, which are stored in a single Feather file for each pulsar.
+The timing residuals are computed using PINT for the NG TOAs, and Tempo2 for the other TOAs.
+These timing residuals are what remain after subtracting an identical deterministic model for pulse arrival time from all TOAs of all PTAs.
+The Feather files, one per pulsar, are the inputs for all remaining analysis, as shown in Fig.
+9.
+The .par and .tim files are no longer needed.
+Table 5: Parameter name equivalences between PINT and Tempo2.
+Physical meaning PINT Tempo2
+Derivative of projected semi-major axis A1DOT XDOT
+Eccentricity ECC E
+The first time derivative of eccentricity EDOT ECCDOT
+Shapiro delay parameter STIGMA STIG, VARSTIGMA
+We perform a “sanity test”, to check that for a given pulsar, the separate Enterprise objects produced for each PTA have identical astrophysical model parameters.
+Although these parameters were identical in the .par files read by Enterprise, if the resulting objects are queried, the names in those produced with PINT are different than in those objects produced by Tempo2.
+The correspondence is shown in Table 5.
+So, after checking that the parameters have identical values and equivalent names, the Enterprise objects are modified to recognize all of the names.
+This ensures that the design matrix $\mathbf{M}$} contains the correct derivatives in any given column.

### v3

@@
 A central result of our analysis is the posterior distribution of the amplitude $A_{gw}$ and exponent $\gamma_{gw}$ of the putative SGWB energy-density spectrum, modeled as a power law in frequency.
-While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance—assessed via a Bayesian odds ratio and noise-marginalized false-alarm probabilities ($p$-values) for three detection statistics—remains below the conventional $5\sigma$ threshold for a confident detection.
+While these results are consistent with a nonzero SGWB amplitude $A_{gw}$, the statistical significance-assessed via a Bayesian odds ratio and noise-marginalized false-alarm probabilities ($p$-values) for three detection statistics-remains below the conventional $5\sigma$ threshold for a confident detection.
 The inter-pulsar timing-residual correlation, reconstructed as a function of angle $\theta$ between the pulsar lines of sight, matches the Hellings and Downs (HD) prediction.
@@
 We use the same tools and methods as the individual PTAs, as implemented in Enterprise [24] and in the next-generation PTA data-analysis package Discovery [25].
-Both are based on parameterized likelihood models, and employ Markov chain Monte Carlo (MCMC) methods to produce posterior samples (here called “posteriors”) which approximate the inferred posterior probability distributions of the model parameters.
-Detection confidence is assessed using both Bayesian and frequentist methods.
-Note: additional tables and plots may be found in the Supplementary Material (SM).
+Both use parameterized likelihood models and Markov chain Monte Carlo (MCMC) to generate posterior samples (“posteriors”) approximating the parameters’ posterior distributions.
+Note: additional details are provided in End Matter (EM) and supporting tables and plots in the Supplementary Material (SM).
 Methods, Models, and Conventions – Because we use standard methods, this paper is brief.
@@
 Again, the timing residuals are assumed to be drawn from a zero-mean Gaussian process, but this time entirely characterized by the covariance matrix ${\bf C_{0}}={\rm diag}\,\bf C$.
+This zeros all elements where the row and column have different pulsars, see A.
 Since our null hypothesis is defined via our signal hypothesis, it has exactly the same parameters/noise-models as the signal hypothesis.
-The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the inter-pulsar HD correlations [28], hence the name “uncorrelated”.
-It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms.
-Thus, CURN does not correspond to the absence of GWs, but rather to “GWs” which create arrival-time fluctuations lacking the inter-pulsar correlations predicted by GR.
-Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2].
-CURN is a more “conservative” null hypothesis than that of “no GWs”.
 Using Discovery [25] as a tool, and conditioning the Gaussian likelihood on the data and the priors, we generate posteriors for the HD and CURN models.
@@
 One result of this analysis is posterior probability distributions for $A_{gw}$ and $\gamma_{gw}$.
-We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to take the value $13/3$.
+We consider models in which $\gamma_{gw}$ is free, and those where it is constrained to $\gamma_{gw}=13/3$.
 The latter would be expected from a population of SMBHBs in circular orbits, provided that GWs dominate the energy loss at nHz frequencies.
 One way to assess detection confidence is via frequentist detection statistics [28].
-We present results for three different statistics that are used within the PTA community.
-All three statistics $D$ are quadratic functions of the timing residuals, with weights determined by $\bf C$ and $\bf C_{0}$.
-Each posterior corresponds to a different model universe, with its own amplitudes/exponents for the different noise sources, and thus its own covariance matrices $\bf C$ and $\bf C_{0}$.
-Hence, the values of the detection statistics $D$, and the probability ($p$-value) of obtaining them under CURN, differ from one posterior sample to the next.
-To assess detection confidence, we present the posterior distributions of these $p$-values.
-In the literature, its mean is called the “posterior predictive $p$-value” [29], similar in spirit to the “noise-marginalized optimal statistic” [30].
-Some past work uses ad hoc methods for assessing the false-alarm probability $p$.
-“Sky scrambling” and “phase shifting” were intended to “remove any signal from the data”.
-The first leaves the data invariant but assigns a large number of random sky positions to the different pulsars, thus destroying the HD correlation [31, 32].
-The second retains the pulsar’s positions, but randomizes the phases of the (frequency-domain Fourier amplitudes of the) data [32].
-However, neither of these methods properly reproduces the null distribution of $D$ [33].
-For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
-We determine the false-alarm probability $p$ using a statistically rigorous method [34, 35, 33, 28].
-Since $D$ is a quadratic form in the data, its distribution under CURN is a generalized $\chi^{2}$ distribution [34, 33, 28]; the probability that it exceeds the observed value in each of our posterior samples is found analytically.
+We consider three different statistics $D$ used within the PTA community.
+Each $D$ is a quadratic function of the timing residuals, with weights determined by $\bf C$ and $\bf C_{0}$.
+Because each posterior sample corresponds to a different noise model and SGWB—and thus different covariances—the resulting statistic values and associated false-alarm probabilities ($p$-values) vary across samples.
+We therefore characterize detection confidence using the posterior distribution of these $p$-values, whose mean corresponds to the posterior predictive $p$-value [29], analogous to the noise-marginalized optimal statistic [30].
+Some past work uses ad hoc techniques to estimate the false-alarm probability $p$, see A.
+We compute $p$-values using a statistically rigorous analytic approach [31, 32, 33, 28].
+Since each statistic $D$ is a quadratic form in the data, its null (CURN) distribution is a generalized $\chi^{2}$; the probability of exceeding the observed value is obtained analytically for each posterior sample.
 To aid interpretation, we also report the $p$-values as “equivalent $q\sigma$” levels: $q>0$ is chosen so that a random draw from a zero-mean Gaussian with variance $\sigma^{2}$ has probability $p$ of exceeding $q\sigma$.
 This “one-sided $p$ value” satisfies $2p=\mathrm{erfc}(q/\sqrt{2})$.
-For example, $1\sigma$ corresponds to $p=16\%$, $2\sigma$ corresponds to $p=2.3\%$, $3\sigma$ corresponds to $p=0.14\%$, $4\sigma$ corresponds to $p=3.2\times 10^{-5}$, and $5\sigma$ corresponds to $p=2.9\times 10^{-7}$.
+For example, $1\sigma$ corresponds to $p=16\%$, $3\sigma$ corresponds to $p=0.14\%$, and $5\sigma$ corresponds to $p=2.9\times 10^{-7}$.
 The latter, a false-alarm probability of less than one per three million, is the traditional threshold for confident detection [12].
@@
 We merge these using the direct combination method [21, 22].
-If not already present, rotation counts are added to the .tim files, which contain pulse TOAs and uncertainties.
-The .par files, containing astrophysical models and detector-specific parameters, are modified one pulsar at a time (the first step of “manual” data combination), as follows.
-  1. 1.
-    For pulsars observed by more than one PTA, a “reference PTA” is selected via the arbitrary ordering $\rm{NG}>\rm{EPTA}>\rm{PPTA}>\rm{MPTA}$.
-InPTA is not listed because its pulsars are observed by other PTAs.
-Other PTAs that have data for this pulsar are called “target PTAs”.
-  2. 2.
-    Since some pulsars have both J- and B-style names, the name is copied from the reference .par file to the target .par files.
-  3. 3.
-    If the pulsar’s .par files use both TCB and TDB time conventions, then quantities whose units contain time are converted to TDB with Tempo2 [41] tools. See SM Sec. III.
-  4. 4.
-    The reference .par file DM model is replaced with a constant-in-time DM model: DM0 is initialized to the .par file DM value at time DMEPOCH, and DM1 and DM2 are initialized to zero.
-  5. 5.
-    A nominal astrophysical model $\beta_{0}$ [1, Eq.
-7.2] (sky position, proper motion, spin frequency, spin-down parameters, binary orbital parameters, and DM values) is copied from the reference .par file to the target .par files.
-Detector-specific parameters in the .par files, for example “FDx” for narrow-band frequency offsets, are unchanged.
-One additional adjustable phase shift parameter (“JUMP”) per target PTA is introduced to account for pulse profile template shape differences.
-The direct combination method [21, 22] adopts one set of parameter values $\beta_{0}$ for each pulsar, item 5 above.
-It works because each pulsar has exact but unknown values $\beta$ for these (for example, for the sky position).
+For each pulsar, this adopts one set of parameter values $\beta_{0}$, defining its nominal astrophysical model [1, Eq.
+7.2]; for details see A and SM Sec.
+A.1.
+It works because each pulsar has exact but unknown values $\beta$ for these parameters (for example, for the sky position).
 They enter the analysis via the nominal value $\beta_{0}$ [1, Eq.
@@
 The design matrix $\mathbf{M}$} allows the effects of DM fluctuations on the timing residuals to be computed from the fixed-in-time DM timing residuals.
-The next analysis step is single-pulsar noise modeling.
-These determine each pulsar’s WN covariance matrix $\mathbf{N}$}, which appears in the Gaussian likelihood [1, Eq.
-7.27].
-This characterizes the timing residual fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eqs.
-7.2 and 7.26], modeled with 30 and 100 frequency bins respectively.
-Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
-Single-pulsar noise modeling determines two or three parameters per pulsar and backend.
-The WN covariance matrix $\mathbf{N}$} is set by these parameters and by the .tim file contents.
-NG, PPTA, MPTA, and seven EPTA backends use the parameters $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and $\mathrm{ECORR}$} [39, Eq.
-2].
-InPTA and the remaining EPTA backends only use $\mathrm{EFAC}$} and $\mathrm{EQUAD}$} [43].
+The next analysis step is single-pulsar noise modeling, which determines each pulsar’s WN covariance matrix $\mathbf{N}$} appearing in the Gaussian likelihood [1, Eq.
+7.27] (see A).
+This modeling yields two or three parameters per pulsar and backend-$\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and (in some cases) $\mathrm{ECORR}$}.
+NG, PPTA, MPTA, and seven EPTA backends use all three parameters [37, Eq.
+2], while InPTA and the remaining EPTA backends use only $\mathrm{EFAC}$} and $\mathrm{EQUAD}$} [40].
 Parameters Uniform on range Stage
@@
 The WN dictionary found in the “S” stage is used in the “F” stage.
-The WN parameters for a given pulsar are obtained from all of its (.tim files) TOA data.
-These are analyzed using JAXopt [44], with priors taken from the Table 1 rows marked “S”; the amplitude/exponent priors are used for all power laws.
+The WN parameters for each pulsar are inferred from all of its .tim-file TOA data using JAXopt [41], with priors from Table 1 (rows marked “S”); the amplitude/exponent priors are applied to all power laws.
 For $\gamma<7$, the $f\to 0$ divergences are absorbed by the timing model, while for $\gamma>1$ the $f\to\infty$ divergences are absent [42, 43].
-Note that $\mathrm{EQUAD}$} has the units used in Tempo [47], Tempo2 [41], and PINT [48], and is written as $\rm T2EQUAD$.
-The output is a “WN dictionary” of best-fit $\mathrm{EFAC}$}, $\mathrm{EQUAD}$}, and (potentially) $\mathrm{ECORR}$} values, used later to obtain $\mathbf{N}$} [39, Eq.
-2].
-For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA analysis values.
-This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more reliable estimates of the WN parameters.
-In SM Sec.
-I, we show that our conclusions are robust: they are not changed much by dropping TOAs from backends with the largest $\mathrm{EFAC}$} corrections.
+The parameter $\mathrm{EQUAD}$} is expressed in the Tempo/Tempo2/PINT convention (T2EQUAD[44, 45, 46].
+The output is a “WN dictionary” of best-fit parameters, from which $\mathbf{N}$} is then determined using the .tim file contents.
 The final stage is multi-pulsar analysis, following the procedures described in [1, 9].
-We use both Enterprise [24] and the next generation PTA data-analysis package, Discovery, which is optimized via JAX for graphics processing units (GPU) [49].
+We use both Enterprise [24] and the next generation PTA data-analysis package, Discovery, which is optimized via JAX for graphics processing units [47].
 The use of Discovery has significantly sped up our analysis.
@@
 1, the other is fixed to $\gamma_{gw}=13/3$.
-Our analysis uses the Hamiltonian No-U-Turn Sampler (NUTS) [50, 51], as implemented in NumPyro [52, 53, 54].
-NUTS is the default for Discovery [25], which provides the log-likelihood gradients needed by Hamiltonian samplers.
-It generates samples that are much more independent than the parallel tempering MCMC sampler [55] previously used by the individual PTAs, and is also much faster.
+Our analysis uses the Discovery [25] default Hamiltonian No-U-Turn Sampler (NUTS) [48, 49], as implemented in NumPyro [50, 51, 52], see A.
 We generated $20{\kern 0.83298pt}480$ posterior samples under the HD signal hypothesis, and the same number under the CURN hypothesis.
@@
 As expected, the posteriors are very similar for the two hypotheses.
-This is because, if there are $\lesssim 1500$ pulsars (see [56, Fig.
+This is because, if there are $\lesssim 1500$ pulsars (see [53, Fig.
 10]), most of the SGWB information is provided by auto- rather than cross-correlations, and by construction, CURN includes these autocorrelations.
-Model $\log_{10}A_{gw}$ $\gamma_{gw}$
-CURN with free $\gamma_{gw}\qquad$ $14.37^{+0.11}_{-0.12}\qquad$ $3.63^{+0.27}_{-0.26}$
-CURN with $\gamma_{gw}=13/3\qquad$ $14.69^{+0.04}_{-0.04}\qquad$
-HD with free $\gamma_{gw}\qquad$ $14.45^{+0.13}_{-0.15}\qquad$ $3.76^{+0.31}_{-0.29}$
-HD with $\gamma_{gw}=13/3\qquad$ $14.72^{+0.04}_{-0.04}\qquad$
-Table 2: Median posterior values for $A_{gw}$ and $\gamma_{gw}$, and their 68% containment intervals (gray bands in Fig. 3).
+Model $\quad\log_{10}A_{gw}$ $\gamma_{gw}$ $\quad\log_{10}\Omega_{gw}$
+CURN ($\gamma_{gw}$ free) $\quad-14.37^{+0.11}_{-0.12}$ $\quad 3.63^{+0.27}_{-0.26}$ $\quad-9.00^{+0.08}_{-0.09}$
+CURN ($\gamma_{gw}$ fixed) $\quad-14.69^{+0.04}_{-0.04}$ $\quad 13/3$ $\quad-8.92^{+0.08}_{-0.08}$
+HD ($\gamma_{gw}$ free) $\quad-14.45^{+0.13}_{-0.15}$ $\quad 3.76^{+0.31}_{-0.29}$ $\quad-9.05^{+0.09}_{-0.09}$
+HD ($\gamma_{gw}$ fixed) $\quad-14.72^{+0.04}_{-0.04}$ $\quad 13/3$ $\quad-8.98^{+0.08}_{-0.09}$
+Table 2: Median posterior values for $A_{gw}$ and $\gamma_{gw}$, their 68% containment intervals from Fig. 2, and $\Omega_{gw}(3\,\text{nHz})$.
 These posteriors, summarized in Table 2, are consistent with the individual PTA results [7, 10, 9, 6].
@@
 2 are smaller (by about a factor of two) than those obtained by intersecting the individual PTA posteriors [54].
-For example, the $\text{area}\equiv\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse for the HD hypothesis is 48% of that shown in [57, Fig.
+For example, the $\text{area}\equiv\int\!{\rm{d}}\gamma_{gw}\int\!{\rm{d}}\log_{10}(A_{gw})$ of our 95.5% containment ellipse for the HD hypothesis is 48% of that shown in [54, Fig.
 1, right].
@@
 Among these, one must show that the data is consistent with the HD signal hypothesis, and very unlikely (at the $5\sigma$ level) to arise under the CURN hypothesis.
-The estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are very similar under both hypotheses.
-One way to assess significance is the frequentist approach: null-hypothesis testing [58] using a statistic.
+The estimates of $A_{GW}$ and $\gamma_{GW}$ cannot help with this, since they are very similar under both hypotheses (see Fig.
+2).
+As previously described, one way to assess significance is via frequentist null-hypothesis testing [55] using a statistic.
 This computes the (noise-marginalized false-alarm) probability $p$ that signal-like data would occur under the null hypothesis.
-We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the data column vector, and consider three choices of the square filter matrix $\mathbf{Q}$} (see [28, Eq.
+We use quadratic statistics of the form $D(z)=z^{\mathsf{T}}{\mathbf{Q}}z$, where $z$ is the column vector of timing residuals, and consider three different standard choices of the square filter matrix $\mathbf{Q}$} (see A and [28, Eq.
 (1)]).
-The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [30, 33, 28].
-The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$ [28].
-The NP minimum-variance (NPMV) statistic [28] adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions.
-Why?
-Normally, the NP statistic would be preferred.
-However, because its filter matrix $\mathbf{Q}$} contains both diagonal (autocorrelation) and off-diagonal (cross-correlation) terms, it is not robust against errors in modeling pulsar noise.
-In contrast, the OS and NPMV statistics use filter matrices that contain only off-diagonal terms, making them more robust.
-Details may be found in [28].
-Figure 3: Histograms of NPMV detection statistic $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
-The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
-The upper horizontal axis is one-sided-Gaussian equivalent significance.
-The detection statistic values and their false-alarm probabilities $p$ depend on the observational data and (inferred) pulsar-noise models under the null (CURN) hypothesis.
-While the data are fixed, the models are uncertain and vary across posteriors, so we adopt a noise-marginalized approach [30, 29].
-For each posterior-sample CURN universe, we compute the three statistics and evaluate their false-alarm probabilities using the corresponding generalized-$\chi^{2}$ distribution [34, 33].
-We then form a histogram of the resulting $p$-values.
-We first discuss the NPMV statistic, whose posterior $p$-value histogram is shown in Fig.
-3.
-The mean/median false-alarm probabilities are $p=5.5\times 10^{-4}$$p=1.1\times 10^{-4}$ for free $\gamma_{gw}$ and $p=3.7\times 10^{-4}$$p=8.0\times 10^{-5}$ for fixed $\gamma_{gw}=13/3$.
-Respectively, these correspond to $3.3\sigma$$3.7\sigma$ and $3.4\sigma$$3.8\sigma$ equivalent Gaussian significances.
-The most likely range is shown in Fig.
-3 as the band that contains 68% of the posterior points about the median.
-These correspond to significances in the range from $3.1\sigma$ to $4.3\sigma$ for free $\gamma_{gw}$ and $3.3\sigma$ to $4.3\sigma$ for $\gamma_{gw}=13/3$.
+While the data are fixed, the inferred pulsar and SGWB noise models-and hence the CURN null-hypothesis distributions-vary across posterior samples.
+We therefore adopt a noise-marginalized approach [30, 29], computing each statistic and its false-alarm probability for every posterior realization using the corresponding generalized $\chi^{2}$ distribution [31, 33].
+The resulting $p$-value distributions are shown in SM Sec.
+A.2.
  statistic  significance mean median 68% range 95% range
@@
 Ranges are centered about the median, and $p$-values are also translated into “equivalent-$\sigma$” units.
-The corresponding histograms for the NP and OS false-alarm probabilities $p$ are similar (see SM Sec.
-II).
-We list key results for all three statistics in Table 3.
+We summarize the results for all three statistics in Table 3.
 None has a mean false-alarm probability $p$ small enough to exceed the $5\sigma$ detection threshold [12].
@@
 Another way to quantify detection confidence [12] is via the Bayes factor ${\mathrm{BF}}=E_{\rm{HD}}/E_{\rm{CURN}}$, the ratio of HD to CURN model evidences.
-Here, the evidence $E$ is the likelihood marginalized over the prior [59]; $E$ has units $({\rm time})^{-N}$ where $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ is the number of TOAs (see SM Sec.
-I Table 4).
-Evaluating these integrals is challenging because the parameter space has $4\times 121+2=486$ dimensions for $\gamma_{gw}$ free ($485$ for $\gamma_{gw}=13/3$).
-To compute the evidences we use generalized stepping-stone sampling [60, 61, 62], which broadens the sharply peaked likelihood using $K=8$ chains (K1=7K-1=7 steps) while restricting the integration region.
-This gives $\ln(E_{\rm{HD}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}504.39\pm 0.10$ and $\ln(E_{\rm{CURN}}\,\rm{s}^{N})=12{\kern 0.83298pt}500{\kern 0.83298pt}494.21\pm 0.09$, implying $\ln({\mathrm{BF}})=10.18\pm 0.13$ and ${\mathrm{BF}}=26{\kern 0.83298pt}000\pm 3{\kern 0.83298pt}000$.
-This is two orders-of-magnitude larger than the Bayes factor ${\mathrm{BF}}\approx 200$ reported by NG [9].
+Evaluating these (see A) gives $\ln({\mathrm{BF}})=10.18\pm 0.13$ $\implies$ ${\mathrm{BF}}=26{\kern 0.83298pt}000\pm 3{\kern 0.83298pt}000$.
+This is two orders-of-magnitude larger than the ${\mathrm{BF}}\approx 200$ reported by NG [9].
 For a real Gaussian random variable, the corresponding significance [29, Eq.
@@
 An optimal method for “reconstructing” the HD correlation from the data has been derived [56] but is not yet fully implemented [43].
-So, our reconstruction is based on the same method [56] previously used by EPTA [7] and NG [9], as implemented in [64].
+So, our reconstruction is based on the same method [53] previously used by EPTA [7] and NG [9], as implemented in [57] (for further details, see A).
 Fig.
 3 shows the result for the $20{\kern 0.83298pt}480$ free $\gamma_{gw}$ CURN posterior samples.
-Our HD reconstruction uses the NG [9] choice of 15 angular bins.
-The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin has 484 pairs.
-For each posterior sample, the correlations of the pulsar pairs are weighted as in [64], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ per bin.
-For each bin, we then construct and average the $20{\kern 0.83298pt}480$ corresponding Gaussian probability distributions.
-The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked.
-Fig.
-4 shows their means (close to medians) and standard deviations.
 A traditional $\chi^{2}$ goodness-of-fit test shows that the reconstruction matches the HD curve well: the reduced $\chi^{2}$ of $0.74$ is consistent with expected statistical fluctuations for 15 degrees of freedom
-Conclusion — We search for an SGWB using public data from five PTAs.
-In effect, we repeat an established analysis, but with a larger dataset.
-This is enabled by the physically motivated step of adopting consistent astrophysical parameters and noise models for all pulsars.
+Conclusion - We search for an SGWB using public data from five PTAs.
+In effect, we repeat an established analysis, but with a larger dataset (see A).
+This is enabled by the physically motivated step of adopting a single astrophysical/noise model for each pulsar.
 Our results provide the strongest evidence to date that an SGWB contributes to observed pulsar pulse-arrival-time fluctuations.
-However, the statistical significance—–assessed using multiple frequentist and Bayesian methods—–remains too low to claim a detection.
+However, the statistical significance—assessed using multiple frequentist and Bayesian methods—remains too low to claim a detection.
 Using existing analysis pipelines enables straightforward comparison with individual PTA results, with which our findings are consistent, but it has limitations.
 For example, these pipelines assume that for a stationary process, the Fourier amplitudes in different frequency bins are statistically independent.
-Although this assumption is not strictly valid [63, see Conclusion], its impact may be negligible or correctable [45].
+Although this assumption is not strictly valid [56, see Conclusion], its impact may be negligible or correctable [42].
 Likewise, our HD-curve reconstruction is suboptimal because implementation of an optimal method [56] is still in progress [43].
@@
 Thus, the reliability of our conclusions depends on the correctness of the five existing PTA analyses.
-We deliberately kept this analysis straightforward and did not “iterate” it.
-This departs from standard pulsar-astronomy practice, where models (e.g., for pulsar timing) are iteratively revised.
-Here, we avoid iterative refinement and tuning of the full PTA search pipeline.
-Past experience with weak-signal detection in LIGO, as well as theoretical studies for PTAs [11], suggests that such iteration can be misleading.
-Because they are not public, we could not include CPTA data [8] in our analysis.
-These are from the Five-hundred-meter Aperture Spherical Telescope (FAST) in Guizhou, China, which is the world’s most sensitive radio telescope.
-Including these high-quality FAST data in future analyses would be interesting and informative.
 Notes: (1) Our full analysis code will be released publicly upon publication.
 (2) After this work was circulated, a different data-combination method was proposed [60].
-Unlike ours, this does not use a physically consistent timing model for a given pulsar.
+Unlike ours, that does not use a single consistent astrophysical/noise model for each pulsar.
 See also SM Sec.
-I.
+A.3.
 Acknowledgments – We thank R.
@@
 Vallisneri for helpful feedback.
+Appendix A End Matter
+EM1: The diagonal operation sets to zero all elements where the row and column have different pulsars, eliminating the inter-pulsar HD correlations [28], hence the “uncorrelated” in CURN.
+It does not remove the dependence upon $A_{gw}$ and $\gamma_{gw}$, which also enter via the diagonal terms.
+Thus, CURN does not correspond to the absence of GWs, but rather to “GWs” which create arrival-time fluctuations lacking the inter-pulsar correlations predicted by GR.
+Alternatively, CURN corresponds to an astrophysical or nuclear/QCD physics process of unknown origin, generating Gaussian rotational fluctuations which are uncorrelated between pulsars, but have the exact same power-law spectrum for all pulsars [2].
+CURN is a more “conservative” null hypothesis than that of “no GWs”.
+EM2: Some previous work determined $p$ values using “sky scrambling” and “phase shifting”, which were intended to “remove any signal from the data”.
+The first leaves the data invariant but assigns a large number of random sky positions to the different pulsars, thus destroying the HD correlation [61, 62].
+The second retains the pulsar’s positions, but randomizes the phases of the (frequency-domain Fourier amplitudes of the) data [62].
+However, neither of these methods properly reproduces the null distribution of $D$ [33].
+For example, because phase shifting does not modify the squared modulus of the Fourier amplitudes, it introduces less variance in $D$ than would arise under the null hypothesis.
+EM3: The data are merged as follows (also see SM Sec.
+A.1).
+If not already present, rotation counts are added to the .tim files, which contain pulse TOAs and uncertainties.
+The .par files, containing astrophysical models and detector-specific parameters, are modified one pulsar at a time (the first step of “manual” data combination), as follows.
+  1. 1.
+    For pulsars observed by more than one PTA, a “reference PTA” is selected via the arbitrary ordering $\rm{NG}>\rm{EPTA}>\rm{PPTA}>\rm{MPTA}$.
+InPTA is not listed because its pulsars are observed by other PTAs.
+Other PTAs that have data for this pulsar are called “target PTAs”.
+  2. 2.
+    Since some pulsars have both J- and B-style names, the name is copied from the reference .par file to the target .par files.
+  3. 3.
+    If the pulsar’s .par files use both TCB and TDB time conventions, then quantities whose units contain time are converted to TDB with Tempo2 [45] tools. See SM Sec. A.1.
+  4. 4.
+    The reference .par file DM model is replaced with a constant-in-time DM model: DM0 is initialized to the .par file DM value at time DMEPOCH, and DM1 and DM2 are initialized to zero.
+  5. 5.
+    A nominal astrophysical model $\beta_{0}$ [1, Eq.
+7.2] (sky position, proper motion, spin frequency, spin-down parameters, binary orbital parameters, and DM values) is copied from the reference .par file to the target .par files.
+Detector-specific parameters in the .par files, for example “FDx” for narrow-band frequency offsets, are unchanged.
+One additional adjustable phase shift parameter (“JUMP”) per target PTA is introduced to account for pulse profile template shape differences.
+EM4: The WN covariance matrix $\mathbf{N}$} characterizes the timing residual fluctuations which remain after subtracting the deterministic timing model determined by $\beta$, and stationary power-law-spectrum models for the RN and DM noise [1, Eqs.
+7.2 and 7.26], modeled with 30 and 100 frequency bins respectively.
+Note that the WN model is nonstationary: for a given pulsar/backend/PTA, the WN variance differs between observations.
+For pulsars with data from multiple PTAs, the WN parameters are shifted away from their single-PTA analysis values.
+This is because the combined data allow better modeling of the low-frequency RN and DM noise, leading to more precise estimates of the WN parameters.
+In SM Sec.
+A.3, we show that our conclusions are robust: they are not changed much by dropping TOAs from backends with the largest $\mathrm{EFAC}$} corrections.
+EM5: NUTS is the default sampler for Discovery [25], which provides the log-likelihood gradients needed by Hamiltonian samplers.
+It generates samples that are much more independent than the parallel tempering MCMC sampler [63] previously used by the individual PTAs, and is also much faster.
+EM6: The optimal statistic (OS) selects $\mathbf{Q}$} to maximize the signal-to-noise ratio (SNR) [30, 33, 28].
+The Neyman–Pearson (NP) statistic selects $\mathbf{Q}$} to maximize the detection probability at fixed false-alarm probability $p$ [28].
+The NP minimum-variance (NPMV) statistic [28] adopts the NP matrix but sets all diagonal entries of $\mathbf{Q}$} to zero, eliminating autocorrelation contributions, making it less sensitive to errors in modeling pulsar noise [28].
+EM7: Here, the evidence $E$ is the likelihood marginalized over the prior [64]; $E$ has units $({\rm time})^{-N}$ where $N=1{\kern 0.83298pt}090{\kern 0.83298pt}206$ is the number of TOAs (see SM Sec.
+A.3 Table 5).
+Evaluating the evidence integrals is challenging because the parameter space has $4\times 121+2=486$ dimensions for $\gamma_{gw}$ free ($485$ for $\gamma_{gw}=13/3$).
+To compute the evidences we use generalized stepping-stone sampling [65, 66, 67], which broadens the sharply peaked likelihood using $K=8$ chains (K1=7K-1=7 steps) while restricting the integration region.
+This gives $\ln(\rm{s}^{N}\,E_{\rm{HD}})=12{\kern 0.83298pt}500{\kern 0.83298pt}504.39\pm 0.10$ and $\ln(\rm{s}^{N}\,E_{\rm{CURN}})=12{\kern 0.83298pt}500{\kern 0.83298pt}494.21\pm 0.09$.
+EM8: Our HD reconstruction uses the NG [9] choice of 15 angular bins.
+The $121\times 120/2=7{\kern 0.83298pt}260$ pulsar pairs are divided equally: each bin has 484 pairs.
+For each posterior sample, the correlations of the pulsar pairs are weighted as in [57], giving $20{\kern 0.83298pt}480$ mean correlations $\mu$ and their variances $\sigma^{2}_{\mu}$ per bin.
+For each bin, we then construct and average the $20{\kern 0.83298pt}480$ corresponding Gaussian probability distributions.
+The resulting 15 (posterior estimators of the) probability distributions of the correlation are sharply peaked.
+Fig.
+3 shows their means (close to medians) and standard deviations.
+EM9: We deliberately kept this analysis straightforward and did not “iterate” it.
+This departs from standard pulsar-astronomy practice, where models (e.g., for pulsar timing) are iteratively revised.
+Here, we avoid iterative refinement and tuning of the full PTA search pipeline.
+Past experience with weak-signal detection in LIGO, as well as theoretical studies for PTAs [11], suggests that such iteration can be misleading.
+Because they are not public, we could not include CPTA data [8] in our analysis.
+These are from the Five-hundred-meter Aperture Spherical Telescope (FAST) in Guizhou, China, which is the world’s most sensitive radio telescope.
+Including these high-quality FAST data in future analyses would be interesting and informative.
 Supplementary Material
-Here, we provide additional plots and tables as supporting material.
-I Effects discarding “outlier backend” TOAs
-In this work, we analyze the full five PTA dataset, which has $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ TOAs.
-We were curious to see if the results would change if we discarded “outlier” data.
-By this, we mean data for which the single pulsar analysis gives WN parameters differing significantly from those naively implied by the .tim file contents.
-We flag such outliers via $\mathrm{EFAC}$}, whose nominal value is (close to) unity, as shown in Fig.
-5.
-Figure 5: Histograms showing how many backends and TOAs have a given value of $\mathrm{EFAC}$}.
-To see the effects of discarding outliers, we drop TOAs from backends where $\mathrm{EFAC}$} lies outside the range $[1/Q,Q]$ for $Q=10,4,3$ and $2$.
-The number of backends/TOAs and fraction of data eliminated by these cuts are shown in Table 4.
-Table 4: Summary of EFAC cuts and their impact on backends and TOAs.
-EFAC Backends Backends TOAs TOAs % TOAs
-range discarded retained discarded retained discarded
-$0,\infty)$ 0 976 0 $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ 0
-$[0.10,10]$ 7 PPTA 969 $1{\kern 0.83298pt}915$ $1{\kern 0.83298pt}088{\kern 0.83298pt}291$ 0.18
-$[0.25,4]$ 6 InPTA, 1 NG, 19 PPTA 960 $8{\kern 0.83298pt}226$ $1{\kern 0.83298pt}081{\kern 0.83298pt}980$ 0.75
-$[0.33,3]$ 12 InPTA, 2 NG, 12 PPTA 950 $17{\kern 0.83298pt}002$ $1{\kern 0.83298pt}073{\kern 0.83298pt}204$ 1.56
-$[0.50,2]$ 1 EPTA, 24 InPTA, 4 NG, 18 PPTA 929 $21{\kern 0.83298pt}340$ $1{\kern 0.83298pt}068{\kern 0.83298pt}866$ 1.96
-The impact of dropping these “outlier” TOAs is negligible, as can be seen from the $A_{gw}$ and $\gamma_{gw}$ posteriors shown in Fig.
-6.
-These should be compared with Fig.
-2 in the paper.
-The tighter data cuts do not significantly affect the posteriors or other conclusions.
-We conclude that our “all are welcome” approach to data inclusion has not significantly skewed or affected our results, but note that discarding the “outlier” data systematically increases the inferred amplitude of the SGWB.
-Figure 6: We illustrate the effects of using different $\mathrm{EFAC}$} thresholds to eliminate data from the most poorly modeled backends.
-The black line shows the CURN posterior from Fig.
-2 in the paper.
-The other curves are constructed after dropping TOAs as shown in Table 4, and demonstrate that the CURN posteriors are not significantly affected by the data cuts.
-Note: each curve shows $20{\kern 0.83298pt}480$ posterior samples.
-A recently posted paper [67] comments about “anomalous WN parameters” in a preliminary draft of this work.
-These arose because that earlier analysis contained phase disconnections between some TOAs observed with different backends.
-This was also the reason that 5 pulsars were given “special” treatment there.
-These phase disconnection issues were fixed by adding rotation counts to the input files.
-II False-alarm probability $p$-value distributions
-In the paper, Figure 3 shows a histogram of the posterior values of the false-alarm probability for the NPMV detection statistic.
-Here, in Figures 7 and 8, we provide the corresponding posterior $p$-value histograms for the Neyman-Pearson (NP) statistic and the traditional “optimal statistic” (OS).
-The means, medians, and widths of these distributions are reported in Table 3 of the paper.
-Figure 7: Histograms of NP detection statistic $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
-The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
-The upper horizontal axis is one-sided-Gaussian equivalent significance.
-Figure 8: Histograms of the traditional optimal detection statistic (OS) $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
-The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
-The upper horizontal axis is one-sided-Gaussian equivalent significance.
-III Data Analysis Workflow
+Here, we provide additional plots, figures, and tables as supporting material.
+A.1 Data Analysis Workflow
 Figure 4: A graphical illustration of the data analysis workflow.
@@
 Our data analysis workflow is illustrated in Fig.
-9.
+4.
 It begins at the top left with .par and .tim files coming from the five PTAs’ public data releases.
@@
 The Feather files, one per pulsar, are the inputs for all remaining analysis, as shown in Fig.
-9.
+4.
 The .par and .tim files are no longer needed.
@@
 This ensures that the design matrix $\mathbf{M}$} contains the correct derivatives in any given column.
+A.2 False-alarm probability $p$-value distributions
+Here, we show histograms of the posterior values of the false-alarm probability for three different detection statistics.
+Figure 5: Histograms of NPMV detection statistic $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
+The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
+The upper horizontal axis is one-sided-Gaussian equivalent significance.
+We begin, in Figure 5, with the posterior $p$-value histograms for the NPMV statistic.
+The mean/median false-alarm probabilities are $p=5.5\times 10^{-4}$$p=1.1\times 10^{-4}$ for free $\gamma_{gw}$ and $p=3.7\times 10^{-4}$$p=8.0\times 10^{-5}$ for fixed $\gamma_{gw}=13/3$.
+Respectively, these correspond to $3.3\sigma$$3.7\sigma$ and $3.4\sigma$$3.8\sigma$ equivalent Gaussian significances.
+The most likely range is shown in Fig.
+5 as the band that contains 68% of the posterior points about the median.
+These correspond to significances in the range from $3.1\sigma$ to $4.3\sigma$ for free $\gamma_{gw}$ and $3.3\sigma$ to $4.3\sigma$ for $\gamma_{gw}=13/3$.
+Figures 6 and 7, are the corresponding histograms for the Neyman–Pearson (NP) statistic and the traditional “optimal statistic” (OS).
+The means, medians, and widths of these distributions are reported in Table 3 of the paper.
+Figure 6: Histograms of NP detection statistic $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
+The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
+The upper horizontal axis is one-sided-Gaussian equivalent significance.
+Figure 7: Histograms of the traditional optimal detection statistic (OS) $p$-values for the $20{\kern 0.83298pt}480$ CURN-hypothesis posterior samples.
+The left panel has $\gamma_{gw}$ free, and the right has $\gamma_{gw}=13/3$.
+The upper horizontal axis is one-sided-Gaussian equivalent significance.
+Why do we consider different statistics?
+Normally, the NP statistic would be preferred.
+However, because its filter matrix $\mathbf{Q}$} contains both diagonal (autocorrelation) and off-diagonal (cross-correlation) terms, it is not robust against errors in modeling pulsar noise.
+In contrast, the OS and NPMV statistics use filter matrices that contain only off-diagonal terms, making them less sensitive to errors in modeling of the pulsar noise.
+Details may be found in [28].
+A.3 Effects discarding “outlier backend” TOAs
+In this work, we analyze the full five PTA dataset, which has $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ TOAs.
+We were curious to see if the results would change if we discarded “outlier” data.
+By this, we mean data for which the single pulsar analysis gives WN parameters differing significantly from those naively implied by the .tim file contents.
+We flag such outliers via $\mathrm{EFAC}$}, whose nominal value is (close to) unity, as shown in Fig.
+8.
+Figure 8: Histograms showing how many backends and TOAs have a given value of $\mathrm{EFAC}$}.
+To see the effects of discarding outliers, we drop TOAs from backends where $\mathrm{EFAC}$} lies outside the range $[1/Q,Q]$ for $Q=10,4,3$ and $2$.
+The number of backends/TOAs and fraction of data eliminated by these cuts are shown in Table 5.
+Table 5: Summary of EFAC cuts and their impact on backends and TOAs.
+EFAC Backends Backends TOAs TOAs % TOAs
+range discarded retained discarded retained discarded
+$0,\infty)$ 0 976 0 $1{\kern 0.83298pt}090{\kern 0.83298pt}206$ 0
+$[0.10,10]$ 7 PPTA 969 $1{\kern 0.83298pt}915$ $1{\kern 0.83298pt}088{\kern 0.83298pt}291$ 0.18
+$[0.25,4]$ 6 InPTA, 1 NG, 19 PPTA 960 $8{\kern 0.83298pt}226$ $1{\kern 0.83298pt}081{\kern 0.83298pt}980$ 0.75
+$[0.33,3]$ 12 InPTA, 2 NG, 12 PPTA 950 $17{\kern 0.83298pt}002$ $1{\kern 0.83298pt}073{\kern 0.83298pt}204$ 1.56
+$[0.50,2]$ 1 EPTA, 24 InPTA, 4 NG, 18 PPTA 929 $21{\kern 0.83298pt}340$ $1{\kern 0.83298pt}068{\kern 0.83298pt}866$ 1.96
+The impact of dropping these “outlier” TOAs is negligible, as can be seen from the $A_{gw}$ and $\gamma_{gw}$ posteriors shown in Fig.
+9.
+These should be compared with Fig.
+2 in the paper.
+The tighter data cuts do not significantly affect the posteriors or other conclusions.
+We conclude that our “all are welcome” approach to data inclusion has not significantly skewed or affected our results, but note that discarding the “outlier” data systematically increases the inferred amplitude of the SGWB.
+Figure 9: We illustrate the effects of using different $\mathrm{EFAC}$} thresholds to eliminate data from the most poorly modeled backends.
+The black line shows the CURN posterior from Fig.
+2 in the paper.
+The other curves are constructed after dropping TOAs as shown in Table 5, and demonstrate that the CURN posteriors are not significantly affected by the data cuts.
+Note: each curve shows $20{\kern 0.83298pt}480$ posterior samples.
+A recently posted paper [60] comments about “anomalous WN parameters” in a preliminary draft of this work.
+These arose because that earlier analysis contained phase disconnections between some TOAs observed with different backends.
+This was also the reason that 5 pulsars were given “special” treatment there.
+We fixed these phase disconnection issues by by adding rotation counts to the input files (see A and SM Sec.
+A.1).